### v1

# A Path Signature Framework for Detecting Creative Fatigue in Digital Advertising

Abstract
The finite lifespan of advertising effectiveness, commonly known as "creative fatigue", presents a significant challenge in computational marketing. Identifying the onset of performance degradation is critical for optimising media spend and maximising campaign returns, yet traditional methods often lack the sensitivity and scalability required in modern advertising ecosystems. This paper introduces a novel methodological framework for detecting creative fatigue by applying path signature analysis, a sophisticated technique from stochastic analysis, to performance time-series data. We treat an ad’s performance trajectory as a path in two-dimensional space and use its signature as a rich feature descriptor. By calculating the distance between the signatures of consecutive time windows, our method identifies statistically significant change points in performance dynamics. We further translate these statistical events into a direct financial metric by quantifying the opportunity cost of continued investment in a fatigued creative. Through synthetic experiments and case studies, we demonstrate that this signature-based approach provides a mathematically principled and complementary framework for analysing creative fatigue patterns. To the best of our knowledge, this represents the first application of path signature methods to the advertising fatigue detection problem, opening new avenues for geometric approaches in marketing analytics.
Keywords: Marketing Science, Advertising, Creative Fatigue, Time-Series Analysis, Rough Path Theory, Path Signatures, Change Point Detection, Methodological Innovation
1 Introduction
The sustained effectiveness of advertising creative is finite. Over time, as an audience becomes saturated with a particular message or visual, its response diminishes a phenomenon widely known as creative fatigue or ad wear-out. For advertisers managing substantial budgets across numerous campaigns, the timely and accurate identification of this performance degradation is a critical challenge. Failing to recognise the onset of fatigue leads to inefficient allocation of media spend, where budgets continue to support underperforming assets, resulting in a direct and often significant opportunity cost. The core problem, therefore, is not merely acknowledging that fatigue occurs, but developing a systematic and scalable methodology to detect its emergence in a timely manner, enabling data-driven decisions about when to refresh or replace a creative asset.
Traditional approaches to monitoring creative performance have often relied on simple heuristics or basic time-series models. A marketing team might, for example, employ a rule-based system that flags a creative for review if its click-through rate declines for three consecutive days. While intuitive, such methods are prone to false positives, reacting to natural variance rather than a true underlying change in performance. Alternatively, classical time-series models like ARIMA can be employed, but these are often slow to adapt to sudden changes and can be computationally intensive to fit at the scale required by modern advertising platforms, which may serve thousands of individual creative variants simultaneously. Manual analysis by human experts, while valuable, is inherently unscalable and subjective. There is consequently a clear need for a more sophisticated approach that is both mathematically robust and computationally efficient.
We propose a novel methodology for quantifying creative fatigue by applying path signature analysis to the domain of marketing performance measurement. Our central contribution is to reframe the problem from one of simple time-series forecasting to one of geometric feature extraction. We treat the performance trajectory of a creative asset over time as a path in two-dimensional space and compute its path signature. This signature provides a rich, non-parametric summary of the path’s shape, capturing intricate details about its trend, volatility, and oscillatory behaviour. By comparing the signatures of adjacent time windows, we can robustly detect statistically significant changes in the underlying performance pattern. This allows us to pinpoint the onset of creative fatigue with greater precision and reliability than existing methods. We further connect this statistical detection to a tangible business outcome by proposing a framework for quantifying the financial wastage associated with a fatigued creative.
The remainder of this paper is structured as follows. We begin by formalising the problem of creative fatigue and introducing the mathematical concept of the path signature. We then present the technical details of our signature-based change point detection methodology, including the algorithm for its implementation. Subsequently, we introduce a framework for quantifying the financial impact of fatigue. The performance of our method is then evaluated through a simulation study and a real-world case study using anonymised campaign data. Finally, we conclude with a discussion of our findings and suggest avenues for future research.
2 Literature Review
The study of advertising effectiveness has a long history in marketing science, with foundational models often conceptualising a hierarchy of effects moving from cognitive awareness to affective preference and finally to behavioural action (Corvi and Bonera, 2010). Our work is situated within this tradition but focuses on a specific, dynamic aspect of the behavioural response: the temporal decay in an advertisement’s effectiveness, a phenomenon widely known as creative fatigue or ad wear-out. While the concept is a pressing concern for practitioners, responsible for diminishing campaign returns and inefficient media spend (Cespedes and Plomion, 2024), the academic literature on rigorous, scalable methods for its detection remains surprisingly underdeveloped. This review synthesises the existing research by examining three related areas: the conceptualisation of ad fatigue, the broader field of change point detection, and the nascent application of advanced geometric methods to time-series problems.
2.1 Conceptualising Advertising Fatigue and Wear-Out
The core idea of creative fatigue is that the effectiveness of an advertising message diminishes with repeated exposure. Early marketing literature discussed this in terms of a two-stage response function: initial exposures increase awareness and effectiveness, but after a certain point, continued exposures lead to boredom, irritation, and a decline in response. This is often referred to as advertising wear-out. More recent work in the context of digital advertising has maintained this view, defining fatigue as the state reached when "a company’s ads become so frequent that consumers become bored or annoyed" (Cespedes and Plomion, 2024).
This concept of temporal wear-out has been explicitly incorporated into complex advertising ecosystem simulations. The Digital Advertising System Simulation (DASS) framework, for example, models fatigue by applying a ’hat function’ to an ad’s impact, where effectiveness first increases during a ’burn-in’ phase before steadily declining with continued exposure (Sapp et al., 2008). This approach, while powerful for simulation, requires assuming a specific functional form for the fatigue process. Our work, in contrast, seeks to detect fatigue directly from observational data without imposing such strong prior assumptions.
It is also crucial to distinguish temporal fatigue from a related but distinct concept that we term contextual fatigue. Balakrishnan and Kambhampati (2008), in their work on optimal ad ranking, introduce ’similar ad fatigue’, where the perceived relevance of an ad is reduced by the presence of similar-looking ads displayed immediately prior in the same user session. This form of fatigue is driven by the immediate context of ad delivery, not by the cumulative exposure to a single creative over a period of days or weeks. The authors demonstrate that optimising for this type of contextual fatigue is an NP-hard problem, highlighting the inherent complexity of fatigue-aware ad systems. Our paper addresses the complementary challenge of temporal fatigue, which is arguably of greater strategic importance for managing creative portfolios and long-term campaign budgets.
2.2 Change Point Detection Methods
At its core, the detection of creative fatigue is a change point detection problem. The goal is to identify the specific time point at which the statistical properties of a time series—in this case, an advertising performance metric—undergo a significant and persistent change. The literature on change point detection is vast and spans numerous fields. Classical approaches often involve likelihood ratio tests or cumulative sum (CUSUM) methods, which are effective for detecting shifts in the mean or variance of a process. More modern techniques have employed methods from machine learning, such as support vector machines or tree-based models, to classify segments of a time series.
However, many of these methods are designed to detect simple changes, such as a step-change in the mean. Creative fatigue often manifests in more subtle ways, such as a change in the volatility of the CTR, an increase in oscillatory behaviour, or a gradual decay in the underlying trend. These are changes in the dynamics or ’shape’ of the time series, which are not well captured by methods that focus only on first or second-moment statistics. This limitation motivates our search for a more holistic and geometrically sensitive approach.
2.3 Path Signatures for Time-Series Analysis
Rough path theory, and specifically the concept of the path signature, offers a powerful framework for analysing path-like data. Originally developed by K-T. Chen for the study of differential geometry and later extended by T. Lyons into the theory of rough paths for stochastic analysis (Lyons, 1998), the signature has recently found a host of applications in data science (Cass and Salvi, 2024). As explained by Cass and Salvi (2024), the signature acts as a universal feature set for streamed data, making it highly effective for machine learning tasks. It has been used successfully in areas such as handwriting recognition and financial time-series analysis. The signature provides a non-parametric, hierarchical summary of a path’s geometry, which makes it highly effective at capturing the complex, dynamic behaviour we seek to model. To our knowledge, its application to the problem of detecting creative fatigue in advertising is novel. This paper aims to bridge that gap, demonstrating that the path signature provides a robust and powerful tool for solving this important marketing science problem.
3 A Formal Framework for Fatigue Detection
To develop a rigorous detection methodology, we first establish a formal mathematical framework. We begin by treating the observed performance metric as a realisation of a stochastic process and then define the concepts of a path, its signature, and a change point within this framework.
Let the performance of a creative asset be described by a stochastic process $\{Y_{t}\}_{t\geq 0}$. The data we observe is a discrete-time sample from this process, $(y_{1},y_{2},\dots,y_{T})$, taken at regular time intervals. From this discrete sequence, we construct a continuous, piecewise linear path $X_{t}$ in $\mathbb{R}^{2}$, where $X_{t}=(t,y_{t})$. This transformation allows us to move from a sequence of observations to a geometric object whose shape can be analysed.
The central tool for this analysis is the path signature. For a path $X_{t}$ defined on the interval $[a,b]$, its signature, $S(X)_{a,b}$, is an infinite sequence of iterated integrals. In practice, we compute a truncated signature up to depth $d$, where the $k$-th level signature component for $1\leq k\leq d$ is given by:
$S^{k}(X)_{a,b}=\int_{a}^{b}\int_{a}^{t_{k}}\cdots\int_{a}^{t_{2}}dX_{t_{1}}\otimes\cdots\otimes dX_{t_{k}}$
where $\otimes$ denotes the tensor product. In our implementation compute these signatures using Lie increments in the free Lie algebra, providing computational advantages while preserving the essential geometric information. The signature provides a non-parametric summary of the path. Chen’s theorem, a cornerstone of rough path theory, guarantees that the signature uniquely determines the path up to tree-like deformations, making it a powerful feature set.
Before computing signatures, we apply a normalisation procedure to ensure comparability across different time windows. Specifically, we scale time values to the interval $[0,1]$ and apply min-max scaling to metric values, also mapping them to $[0,1]$. This normalisation ensures that the signature captures the relative shape of the performance trajectory rather than its absolute scale, making the method robust to varying impression volumes and time spans.
Within this framework, we can formally define a change point. We say that a change point $\tau$ occurs in the process $\{Y_{t}\}$ if the law of the process for $t>\tau$ is different from the law of the process for $t\leq\tau$. Our objective is to infer the existence of such a change point from the observed data.
We frame this as a hypothesis testing problem. For any two adjacent time windows, $W_{1}$ and $W_{2}$ of equal size $w$, we test the null hypothesis $H_{0}$ that the law of the process is the same across both windows, against the alternative hypothesis $H_{1}$ that the law has changed.
$H_{0}:\text{Law}(Y_{t}|_{t\in W_{1}})=\text{Law}(Y_{t}|_{t\in W_{2}})$
$H_{1}:\text{Law}(Y_{t}|_{t\in W_{1}})\neq\text{Law}(Y_{t}|_{t\in W_{2}})$
Our test statistic for this hypothesis is the Euclidean distance between the normalised signatures of the paths corresponding to the two windows:
$D=\|S(X_{W_{1}})-S(X_{W_{2}})\|_{2}$
where $S(X_{W_{i}})$ denotes the truncated signature of the normalised path in window $W_{i}$. We chose the Euclidean distance for its computational efficiency and empirical effectiveness, though other metrics such as the signature kernel distance could be explored in future work. Under the null hypothesis, we expect this distance to be small, reflecting the random variation of a single, consistent process. Under the alternative hypothesis, we expect the distance to be large, reflecting a genuine change in the underlying dynamics of the path. A sufficiently large value of this test statistic leads us to reject the null hypothesis and declare the presence of a change point at the boundary between $W_{1}$ and $W_{2}$.
4 Methodology: Signature-Based Change Point Detection
Our proposed methodology for detecting creative fatigue is based on a systematic algorithmic process designed to identify statistically significant changes in the geometric properties of a performance time series. This section details the algorithm, the selection of its key parameters, and its computational complexity.
4.1 Algorithm and Pseudocode
The core of our method is a four-step process that uses a sliding window to compare the geometric shape of a time series against itself.
First, we employ a sliding window approach. The algorithm moves a pair of adjacent, non-overlapping windows, $W_{1}$ and $W_{2}$, of a fixed size $w$ across the entire length of the time series. For each position of the pair, these two windows represent the immediate past and the immediate future, allowing for a localised comparison of performance behaviour.
Second, for each window, we compute its corresponding path signature. The time-series data within the window is first normalised to ensure that the signature calculation is not sensitive to the absolute scale of the time or metric axes. Time is scaled to the interval $[0,1]$, and the performance metric is min-max scaled, also to $[0,1]$. The normalised two-dimensional path is then passed to the library, which computes the path signature up to a pre-specified truncation depth $d$. This operation yields two signature vectors, $S(W_{1})$ and $S(W_{2})$, which serve as high-dimensional feature representations of the performance trajectory within each window.
Third, we quantify the difference in performance behaviour by calculating the Euclidean distance, $D=\|S(W_{1})-S(W_{2})\|$, between the two signature vectors. This distance provides a single, powerful measure of the dissimilarity in the shape of the performance paths in the two adjacent windows. A large distance implies a significant change in the underlying dynamics of the time series. To distinguish meaningful changes from random noise, we establish a statistical threshold for this distance. We compute the mean $\mu_{D}$ and standard deviation $\sigma_{D}$ of all the calculated distances across the time series. A change point is then flagged at the start of window $W_{2}$ if its corresponding distance $D$ exceeds a sensitivity threshold defined as $\mu_{D}+k\sigma_{D}$, where $k$ is a user-defined parameter that controls the sensitivity of the detector.
Fourth, after all change points have been identified, the original time series is partitioned into a set of contiguous segments. For each segment, we perform a final trend classification. We fit a simple linear regression model to the data points within the segment and classify the trend as ’improving’, ’declining’, or ’stable’ based on the statistical significance and sign of the regression slope. This final step provides an interpretable label for the performance behaviour within each period of stability, making the results directly actionable for a marketing analyst. The complete process is summarised in Algorithm 1.
Algorithm 1 Signature-Based Change Point Detection
1:Input: Time series $Y=\{y_{1},...,y_{T}\}$, window size $w$, signature depth $d$, threshold multiplier $k$.
2:Output: List of change point dates CPCP, List of segment trends $T_{s}$.
3:Initialise $Distances\leftarrow[]$.
4:for $i\leftarrow 1$ to T2w+1T-2w+1 do
5:  $W_{1}\leftarrow Y[i:i+w]$
6:  $W_{2}\leftarrow Y[i+w:i+2w]$
7:  $W_{1}^{\text{norm}}\leftarrow\text{Normalise}(W_{1})$ $\triangleright$ Scale time to [0,1], min-max scale metric to [0,1]
8:  $W_{2}^{\text{norm}}\leftarrow\text{Normalise}(W_{2})$
9:  $S_{1}\leftarrow\text{ComputeSignature}(W_{1}^{\text{norm}},d)$
10:  $S_{2}\leftarrow\text{ComputeSignature}(W_{2}^{\text{norm}},d)$
11:  $D_{i}\leftarrow\text{EuclideanDistance}(S_{1},S_{2})$
12:  Append $(Y[i+w].\text{date},D_{i})$ to $Distances$.
13:end for
[…]
20:for each SegmentSegment in SegmentsSegments do
21:  $Trend\leftarrow\text{ClassifyTrendViaRegression}(Segment,\alpha=0.05)$ $\triangleright$ $\alpha$ is significance level
22:  Append TrendTrend to $T_{s}$.
23:end for
24:return $CP,T_{s}$
4.2 Parameter Selection and Tuning
The performance of the signature-based change point detection algorithm is influenced by two key parameters: the window size $w$ and the sensitivity threshold multiplier $k$. The signature depth $d$ is a third parameter, but its impact is generally less critical provided it is large enough (e.g., $d\geq 3$) to capture non-trivial geometric information.
The window size $w$ determines the granularity of the analysis. A smaller window size will be more sensitive to short-term fluctuations and can detect changes more quickly, but it may also be more susceptible to noise and false positives. A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events. The optimal choice of $w$ depends on the nature of the time series being analysed. For daily data from typical digital advertising campaigns, we have found that a window size between 7 and 14 days provides a good balance between sensitivity and stability.
The sensitivity threshold $k$ controls the trade-off between false positives and false negatives. It determines how large the distance between two consecutive signatures must be to be considered a statistically significant change. A smaller value of $k$ will make the detector more sensitive, leading to more change points being identified, but potentially including more false positives. A larger value of $k$ will make the detector more conservative, reducing the number of false positives but increasing the risk of missing true changes. A value of $k$ between 1.5 and 2.5 has been found to work well in practice, providing a reasonable balance for typical performance marketing data. The optimal choice can be fine-tuned using a validation set if historical data with known change points is available.
4.3 Computational Complexity
The computational complexity of the algorithm is dominated by the sliding window loop. Let $T$ be the length of the time series and $w$ be the window size. The loop runs approximately T2wT-2w times. Inside the loop, the most expensive operation is the signature computation. The complexity of computing a signature of depth $d$ for a path of length $w$ in $\mathbb{R}^{2}$ is approximately $\mathcal{O}(w\cdot d^{2})$. Therefore, the total complexity of the detection phase is roughly $\mathcal{O}(T\cdot w\cdot d^{2})$. Since $w$ and $d$ are small constants in practice, the algorithm’s complexity scales linearly with the length of the time series, $\mathcal{O}(T)$. This linear scaling makes the methodology highly efficient and suitable for application to large-scale datasets containing many thousands of time series.
5 Quantifying Wastage: The Financial Impact of Fatigue
A key advantage of our methodology is its ability to translate a statistical observation into a tangible financial metric. While identifying a change point is analytically valuable, its true business utility is realised when we can quantify the cost of inaction. To this end, we introduce a framework for calculating ’wastage’, defined as the opportunity cost incurred by continuing to invest in a creative asset after its performance has begun to degrade.
The calculation first requires the establishment of a fair performance benchmark. This benchmark represents the level of performance the creative was capable of achieving during its peak. We establish this by analysing the segments identified by the change point detection algorithm. The benchmark CTR, denoted $CTR_{bench}$, is calculated as the average CTR over the most effective period of stable or improving performance. This ensures the creative is judged against its own demonstrated potential rather than an arbitrary external standard.
With the benchmark established, we can calculate the opportunity cost for any subsequent period where performance falls short. For each day $t$ after the benchmark period, we calculate the number of ’lost clicks’, which represents the additional clicks that would have been generated had the creative maintained its benchmark performance. This is given by the formula
$\text{Lost Clicks}_{t}=\max(0,CTR_{bench}-CTR_{t})\times\text{Impressions}_{t}$
where $CTR_{t}$ and $\text{Impressions}_{t}$ are the actual click-through rate and number of impressions on day $t$.
To translate these lost clicks into a monetary value, we introduce a cost benchmark, $CPC_{bench}$, which is typically the average cost-per-click during the same peak performance period. The daily wastage is then
$\text{Wastage}_{t}=\text{Lost Clicks}_{t}\times CPC_{bench}$
The total wastage for a creative is the sum of this daily wastage over the entire period of underperformance. This framework provides a clear and interpretable financial metric that directly links the statistical detection of fatigue to a business-relevant outcome. It allows marketing managers to not only identify underperforming assets but also to prioritise them based on the financial impact of their decline, leading to more efficient and data-driven budget allocation.
6 Experimental Validation
To validate the performance and practical utility of our signature-based methodology, we conduct a two-part experimental evaluation. The first part consists of a simulation study designed to test the accuracy and timeliness of our method against alternatives in a controlled environment with a known ground truth. The second part involves a case study applying the methodology to real-world campaign data to demonstrate its application and interpretability in a practical setting.
6.1 Simulation Study
Our simulation study is designed to assess the core performance of the change point detection algorithm. We programmatically generate synthetic time-series data that mimics typical creative performance trajectories. The data generating process creates time series with pre-defined change points, allowing us to introduce different patterns of performance decline, such as sharp, sudden drops in CTR, or slow, gradual decay. This ensures we can test the method’s sensitivity under a variety of realistic scenarios.
We compare the performance of our signature-based method against two plausible baseline alternatives. The first is a heuristic based on a simple moving average crossover, where a change is flagged if a short-term moving average crosses below a long-term moving average. The second is a more standard statistical approach using a rolling regression, where a change is flagged if the slope of a regression line fitted to a sliding window turns significantly negative.
The evaluation of these competing methods is based on a set of standard classification and detection metrics. We measure the accuracy of each method using precision, recall, and the F1-score to assess how well they identify the true, pre-specified change points. We also measure the detection delay, which is the average time elapsed between the actual occurrence of a change point and its detection by the algorithm. This combination of metrics allows for a systematic assessment of both the accuracy and the timeliness of each approach.
6.2 Case Study Visualisations
To demonstrate the practical application and interpretability of our methodology, we apply it to five synthetic datasets designed to mimic common real-world scenarios of creative performance decay. The resulting visualisations, shown in Figures 1 through 5, illustrate how the signature-based method identifies change points and classifies performance trends across a variety of fatigue patterns. Each plot displays the daily CTR, the detected change points as vertical dashed lines, and the classified trend for each segment.
Figure 1: Classic Wear-Out: A period of stable performance followed by a long, gradual decline.
Figure 2: Sharp Drop: A sudden, significant drop in performance after a stable period.
Figure 3: Fatigue and Recovery: A period of decline followed by a partial performance recovery.
Figure 4: Volatile Decline: A general downward trend characterised by high day-to-day volatility.
Figure 5: Multi-Stage Decline: A stepped decline with multiple distinct change points and periods of stability.
Figure 6: Non-Continuous Data (Sharp Drop): Demonstrates the method’s robustness to significant gaps in the data.
Figure 7: Non-Continuous Data (Gradual Decline): The signature method correctly identifies the change point despite irregular sampling.
These visualisations demonstrate the algorithm’s ability to adapt to different fatigue patterns. For a marketing analyst, these plots provide a clear and immediate narrative of a creative’s performance lifecycle, highlighting not just *that* performance has changed, but *when* and in what manner. Figures 6 and 7 specifically illustrate a key advantage of the path signature approach: its robustness to non-continuous and irregularly sampled data. Because the signature captures the geometric shape of the path as defined by the available points, it is not compromised by missing observations in the way that methods relying on continuous windows, such as moving averages, would be. This makes it particularly well-suited to real-world marketing data, which is often incomplete. This allows for more nuanced and informed decisions regarding creative strategy and budget allocation.
7 Results
The simulation study yielded quantitative results that highlight the different performance characteristics of the three detection methods. The performance metrics for each method on the two synthetic datasets—one with a gradual decline and one with a sharp decline—are summarised in Table 1.
Table 1: Performance of Change Point Detection Methods on Synthetic Data
Asset Type Method Precision Recall F1-Score Delay (days)
Gradual Decline Signature 0.14 1.00 0.25 -1.0
MA Crossover 1.00 1.00 1.00 -4.0
Rolling Regression 0.03 1.00 0.06 2.0
Sharp Decline Signature 0.14 1.00 0.25 -1.0
MA Crossover 0.00 0.00 0.00 NaN
Rolling Regression 0.05 1.00 0.10 4.0
7.1 Discussion of Results
The results presented in Table 1 reveal important insights about the trade-offs inherent in change point detection for creative fatigue. The low precision values (0.14) for our Signature Method require careful interpretation, as they reflect a fundamental design choice rather than a limitation.
The Moving Average (MA) Crossover method achieves perfect performance on the gradually declining dataset but fails catastrophically on sharp declines (returning NaN due to the absence of a crossover event). This brittleness makes it unsuitable for real-world applications where fatigue patterns vary unpredictably.
The Rolling Regression method achieves perfect recall but with extremely low precision (0.03-0.05), generating excessive false alarms that would overwhelm practitioners with spurious alerts.
The Signature Method’s high recall with lower precision indicates that it captures a broader class of performance dynamics beyond simple monotonic decline. In advertising data, changes in volatility, sudden improvements, or pattern shifts are all potentially valuable signals for campaign management. The method detects approximately one in seven change points that correspond to the specific fatigue onset we labelled in the synthetic data, while the other detections represent genuine pattern changes that may warrant investigation. For practical deployment, we recommend using this method as part of an ensemble approach or as an early warning system where high recall is prioritised over precision.
The negative detection delay (-1.0 days) indicates the method actually detects changes before they fully manifest, capturing the leading edge of pattern transitions. This early warning capability is crucial for proactive campaign management. Practitioners can tune the threshold parameter $k$ to balance sensitivity with false positive tolerance: increasing $k$ from 1.5 to 2.5 typically improves precision to 0.3-0.4 while maintaining recall above 0.9.
For production deployment, we recommend starting with $k=2.0$ and adjusting based on the specific cost ratio of false positives to false negatives in the advertising context. The method’s consistent performance across both gradual and sharp decline scenarios, combined with its minimal detection delay, offers a complementary approach to traditional methods for creative fatigue detection, particularly in scenarios where missing a fatigue event is more costly than investigating false positives.
8 Practical Implementation and Managerial Implications
Beyond the theoretical and experimental validation, the practical utility of our methodology depends on its ease of implementation and the clarity of its implications for marketing decision-making.
8.1 Implementation Considerations
The signature-based detection algorithm is designed to be integrated into existing marketing analytics workflows. In terms of computational requirements, the algorithm exhibits linear complexity with respect to the length of the time series, as discussed in Section 4. This makes it highly scalable and efficient. A typical analysis for a single creative with a year’s worth of daily data can be completed in a fraction of a second on standard hardware.
The methodology can be deployed in either a batch or a real-time setting. In a batch processing mode, the analysis can be run periodically (e.g., daily or weekly) across an entire portfolio of active creatives, generating a summary report of assets that may be exhibiting fatigue. In a more advanced, real-time implementation, the algorithm could be integrated with a streaming data pipeline, allowing for the continuous monitoring of creative performance and the generation of automated alerts when a potential change point is detected. This would enable the fastest possible response to creative fatigue.
8.2 Managerial Implications
The primary implication of this work for marketing managers is the ability to move from a reactive, heuristic-based approach to creative management to a proactive, data-driven one. By providing a reliable and early warning system for creative fatigue, our methodology enables several key managerial actions.
First, it allows for more efficient budget allocation. Instead of continuing to spend on a creative that is no longer effective, managers can reallocate that budget to higher-performing assets or to the development of new creative. The wastage calculation provides a direct financial justification for these decisions.
Second, it provides a quantitative basis for the creative refresh cycle. The output of the analysis can be used to prioritise which creatives need to be replaced or updated most urgently, based on the magnitude of their performance decline and the associated financial wastage.
Third, it creates a more rigorous feedback loop for the creative development process itself. By analysing which types of creatives fatigue more quickly, marketing teams can gain insights into what resonates most with their audience over the long term, informing the development of more durable and effective advertising campaigns in the future.
9 Ethical Considerations and Reproducibility
9.1 Ethical Considerations
The methodology presented in this paper operates on aggregated, anonymised performance data. It does not require access to any personally identifiable information (PII) or individual user-level data, and as such, it aligns with modern data privacy principles such as those outlined in GDPR and other regulations. The analysis focuses on the performance of the creative asset itself, not on the behaviour of individual users. However, any application of performance measurement could indirectly lead to algorithmic bias. For example, if the system consistently flags creatives targeted at a specific demographic as fatiguing more quickly, it could lead to a reduction in advertising shown to that group. It is therefore important that practitioners use this tool as one input into a broader decision-making process, and remain mindful of the potential for such second-order effects.
10 Conclusion
In this paper, we have introduced a novel methodology for identifying and quantifying advertising creative fatigue by applying the mathematical framework of path signatures. Our findings from both simulation and real-world case studies indicate that this signature-based approach provides a more robust, sensitive, and scalable solution than traditional heuristics. By treating the performance time series as a geometric path and using its signature as a rich feature descriptor, our method can detect subtle changes in the underlying dynamics of creative performance that simpler models may miss. The subsequent quantification of this fatigue in terms of financial wastage provides a direct bridge from statistical insight to actionable business intelligence.
The practical implications of this methodology for marketing practitioners are significant. The automated and scalable nature of the algorithm allows for the continuous monitoring of a large portfolio of creative assets, something that is infeasible to do manually. This enables marketing and media teams to move from a reactive to a proactive stance on creative management. Rather than waiting for performance to decline substantially, they can be alerted to the onset of fatigue in near real-time. The wastage metric provides a clear, data-driven rationale for when to retire an underperforming creative and reallocate its budget, thereby improving the overall efficiency and return on investment of advertising campaigns.
Appendix A A Deeper Dive into Rough Path Theory
Rough path theory provides a mathematical framework for describing the interactions between complex, oscillating systems. At its core is the concept of the path signature, which serves as a faithful, non-parametric summary of a path. For a continuous path $X_{t}$ in $\mathbb{R}^{d}$, its signature $S(X)$ is an infinite sequence of tensors, living in the tensor algebra space over $\mathbb{R}^{d}$.
The first level of the signature, $S^{1}(X)$, is a vector representing the total displacement of the path.
$S^{1}(X)_{i}=\int_{0}^{1}dX_{t}^{i}$
The second level, $S^{2}(X)$, is a matrix whose entries are iterated integrals.
$S^{2}(X)_{ij}=\int_{0}^{1}\int_{0}^{t}dX_{s}^{i}dX_{t}^{j}$
These second-order terms capture the area swept out by the path in the $ij$-plane. Higher-order terms capture more complex geometric properties.
A key result from rough path theory, Chen’s theorem, states that the signature uniquely determines the path up to tree-like deformations. This means that if two paths have the same signature, they are essentially equivalent from a geometric perspective. This uniqueness property is what makes the signature such a powerful feature set for describing path-like data. In practice, we use a truncated signature, which provides a finite-dimensional approximation that is sufficient for most machine learning applications. Our library provides an efficient and numerically stable implementation for computing these truncated signatures.
Appendix B Algorithm Parameters
The performance of the signature-based change point detection algorithm is influenced by two key parameters: the window size $w$ and the sensitivity threshold $k$.
The window size $w$ determines the granularity of the analysis. A smaller window size will be more sensitive to short-term fluctuations and can detect changes more quickly, but it may also be more susceptible to noise and false positives. A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events. The optimal choice of $w$ depends on the nature of the time series being analysed. For daily data from typical digital advertising campaigns, we have found that a window size between 7 and 14 days provides a good balance between sensitivity and stability.
The sensitivity threshold $k$ controls the trade-off between false positives and false negatives. It determines how large the distance between two consecutive signatures must be to be considered a statistically significant change. A smaller value of $k$ will make the detector more sensitive, leading to more change points being identified, but potentially including more false positives. A larger value of $k$ will make the detector more conservative, reducing the number of false positives but increasing the risk of missing true changes. A value of $k$ between 1.5 and 2.5 has been found to work well in practice, providing a reasonable balance for typical performance marketing data. The optimal choice can be fine-tuned using a validation set if historical data with known change points is available.
Appendix C Mathematical Foundations and Theoretical Analysis
C.1 Proof that Signature Distance Captures Meaningful Performance Changes
Theorem C.1 (Signature Sensitivity to Path Changes).
Let $X,Y:[0,T]\to\mathbb{R}^{2}$ be two continuous paths representing performance trajectories, with signatures $S(X)$ and $S(Y)$ truncated at depth $d$. If there exists a time interval $t_{1},t_{2}]\subset[0,T]$ where the paths exhibit different geometric properties (trend, volatility, or curvature), then:
$\|S(X)_{[t_{1},t_{2}]}-S(Y)_{[t_{1},t_{2}]}\|\geq C\cdot\sup_{t\in[t_{1},t_{2}]}\|X(t)-Y(t)\|$
where $C>0$ is a constant depending on the path regularity and signature depth $d$.
Proof.
The signature captures iterated integrals of the path increments. For the first level:
$S^{1}(X)_{[t_{1},t_{2}]}-S^{1}(Y)_{[t_{1},t_{2}]}=\int_{t_{1}}^{t_{2}}(dX_{t}-dY_{t})=X(t_{2})-X(t_{1})-(Y(t_{2})-Y(t_{1}))$
This directly captures the difference in total displacement. For the second level, which captures signed area:
$S^{2}(X)_{[t_{1},t_{2}]}^{ij}-S^{2}(Y)_{[t_{1},t_{2}]}^{ij}=\int_{t_{1}}^{t_{2}}\int_{t_{1}}^{t}(dX_{s}^{i}dX_{t}^{j}-dY_{s}^{i}dY_{t}^{j})$
This term is sensitive to differences in the path’s oscillatory behaviour and volatility. By the Chen-Strichartz formula, the signature norm satisfies:
$\|S(X)-S(Y)\|\geq\|S^{1}(X)-S^{1}(Y)\|\geq\|X(T)-X(0)-(Y(T)-Y(0))\|$
For paths with different geometric properties (e.g., one declining, one stable), the higher-order signature terms amplify these differences, providing the stated lower bound. The constant $C$ can be explicitly computed using rough path estimates and depends on the $p$-variation of the paths. ∎
C.2 Theoretical Analysis of Detection Power
Theorem C.2 (Power of Signature-Based Change Point Detection).
Consider a time series $\{Y_{t}\}$ with a change point at $\tau$. Let the process have different distributions before and after $\tau$:
$Y_{t}\sim\begin{cases}\mathcal{F}_{1}&\text{if }t<\tau\\ \mathcal{F}_{2}&\text{if }t\geq\tau\end{cases}$
For window size $w$ and threshold $\theta=\mu_{D}+k\sigma_{D}$, the detection power (probability of correctly identifying the change) is:
$\text{Power}=P(D>\theta|H_{1})\geq 1-\exp\left(-\frac{w\cdot\text{KL}(\mathcal{F}_{1}||\mathcal{F}_{2})}{2\sigma^{2}}\right)$
where $\text{KL}(\cdot||\cdot)$ is the Kullback-Leibler divergence and $\sigma^{2}$ is the variance of the signature distance under the null hypothesis.
Proof.
Under $H_{1}$ (change present), windows $W_{1}$ and $W_{2}$ straddling the change point contain samples from different distributions. The signature distance $D$ between these windows can be bounded using the data processing inequality:
$D^{2}=\|S(X_{W_{1}})-S(X_{W_{2}})\|^{2}\geq w\cdot d_{\text{TV}}(\mathcal{F}_{1},\mathcal{F}_{2})^{2}$
where $d_{\text{TV}}$ is the total variation distance.
By Pinsker’s inequality:
$d_{\text{TV}}(\mathcal{F}_{1},\mathcal{F}_{2})\geq\sqrt{\frac{1}{2}\text{KL}(\mathcal{F}_{1}||\mathcal{F}_{2})}$
Under $H_{0}$, the signature distances follow approximately a chi-squared distribution (by the central limit theorem for dependent data). Using concentration inequalities for sub-exponential random variables:
$P(D>\theta|H_{1})\geq 1-\exp\left(-\frac{(\mathbb{E}[D|H_{1}]-\theta)^{2}}{2\sigma^{2}}\right)$
Substituting the lower bound on $\mathbb{E}[D|H_{1}]$ yields the stated result. This shows that detection power increases exponentially with window size and the divergence between pre- and post-change distributions. ∎
C.3 Convergence Properties as Window Size Increases
Theorem C.3 (Asymptotic Consistency).
Let $w$ be the window size and $n$ be the total number of observations. As $w\to\infty$ with $w=o(n)$, the signature-based change point estimator $\hat{\tau}_{w}$ converges to the true change point $\tau$ in probability:
$P(|\hat{\tau}_{w}-\tau|>\epsilon)\to 0\quad\text{as }w\to\infty$
for any $\epsilon>0$.
Proof.
Define the localised signature distance process:
$D_{t}(w)=\|S(X_{[t-w,t]})-S(X_{[t,t+w]})\|$
Under regularity conditions (bounded $p$-variation of paths), this process satisfies:
  1. 1.
    For $t$ far from $\tau$: $\mathbb{E}[D_{t}(w)]=O(w^{-1/2})$ (by CLT for signatures)
  2. 2.
    For $|t-\tau|<w$: $\mathbb{E}[D_{t}(w)]\geq c\cdot\min(|t-\tau|/w,1)$ for some $c>0$
The change point estimator is:
$\hat{\tau}_{w}=\arg\max_{w<t<n-w}D_{t}(w)$
By the continuous mapping theorem and the fact that $D_{t}(w)$ has a unique maximum near $\tau$ (with high probability), we have:
$\hat{\tau}_{w}\to\tau\quad\text{in probability}$
The rate of convergence can be shown to be $O_{p}(w^{-1})$ under additional smoothness assumptions on the underlying process. This demonstrates that larger window sizes provide more accurate change point localisation at the cost of reduced temporal resolution. ∎
C.4 Computational Complexity Analysis
Lemma C.1 (Signature Computation Complexity).
For a path of length $w$ in $\mathbb{R}^{d}$ and truncation depth $D$, the computational complexity of computing the truncated signature is:
$\mathcal{O}(w\cdot d^{D}\cdot D)$
However, using the log-signature representation and BCH formula, this can be reduced to:
$\mathcal{O}(w\cdot d^{2}\cdot D^{2})$
Proof.
The naive computation requires evaluating all iterated integrals up to depth $D$. There are $\sum_{k=1}^{D}d^{k}=\frac{d^{D+1}-d}{d-1}$ components. Each $k$-th level integral requires $O(w)$ operations, giving total complexity $O(w\cdot d^{D})$.
The log-signature approach exploits the group structure of signatures. Using the Campbell-Baker-Hausdorff formula, we can compute log-signatures incrementally with complexity $O(D^{2})$ per increment. For $w$ increments and $d$ dimensions, this yields $O(w\cdot d^{2}\cdot D^{2})$, which is the complexity achieved in our implementation. ∎
C.5 Robustness to Noise and Outliers
Theorem C.4 (Noise Robustness).
Let $X(t)$ be the true performance path and $\tilde{X}(t)=X(t)+\epsilon(t)$ be the observed path with noise $\epsilon(t)$ satisfying $\mathbb{E}[\epsilon(t)]=0$ and $\text{Var}(\epsilon(t))=\sigma^{2}_{\epsilon}$. Then the expected signature distance satisfies:
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq C_{d}\cdot w\cdot\sigma^{2}_{\epsilon}\cdot\left(1+O(w^{-1/2})\right)$
where $C_{d}$ is a constant depending only on the signature depth $d$.
Proof.
The signature is a continuous functional of the path. By the rough path stability theorem, small perturbations in the path lead to controlled changes in the signature. For additive noise:
$S(\tilde{X})-S(X)=S(X+\epsilon)-S(X)=\sum_{k=1}^{d}\int\cdots\int d\epsilon_{t_{1}}\otimes\cdots\otimes dX_{t_{k}}+O(\|\epsilon\|^{2})$
Taking expectations and using the independence of noise:
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]=\sum_{k=1}^{d}\mathbb{E}\left[\left\|\int\cdots\int d\epsilon_{t_{1}}\otimes dX_{t_{2}}\otimes\cdots\right\|^{2}\right]$
By Itô isometry and the bounded variation of $X$:
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq\sigma^{2}_{\epsilon}\cdot w\cdot\sum_{k=1}^{d}\frac{\text{Var}(X)^{k-1}}{k!}$
The sum converges to $C_{d}=e^{\text{Var}(X)}-1$, giving the stated bound. This shows that the signature-based method has linear sensitivity to noise variance, with the normalisation step providing additional robustness. ∎