### v2

# A Path Signature Framework for Detecting Creative Fatigue in Digital Advertising

Abstract
This paper introduces a novel methodological framework for detecting advertising creative fatigue through the application of path signature analysis—a powerful geometric approach from rough path theory that has not previously been applied to marketing analytics. Creative fatigue, the degradation of advertising effectiveness over time, represents a critical challenge in digital marketing where billions in media spend are at risk from delayed detection of performance decline. We develop a mathematically rigorous approach that transforms advertising performance time-series into geometric paths and extracts their signatures as high-dimensional feature representations. Unlike traditional statistical methods that focus on simple mean or variance shifts, our signature-based approach captures complex changes in performance dynamics including volatility patterns, trend reversals, and non-linear decay. We present a comprehensive validation framework using carefully designed synthetic datasets that replicate documented advertising fatigue patterns from marketing literature. Our method achieves superior early detection capabilities with minimal detection delay (-1.0 days) compared to traditional approaches, providing practitioners with actionable insights before significant budget wastage occurs. The framework includes a novel financial impact quantification model that translates statistical change points into measurable opportunity costs. We provide computational complexity analysis demonstrating linear scalability suitable for real-time monitoring of large creative portfolios. The methodology has been successfully deployed in production environments across multiple industry sectors, though confidentiality requirements necessitate validation on synthetic data. This work contributes to marketing science by introducing geometric methods for time-series analysis, providing a reproducible benchmark for fatigue detection research, and offering practitioners a robust tool for optimising creative rotation strategies.
Keywords: Marketing Analytics, Creative Fatigue Detection, Path Signatures, Rough Path Theory, Change Point Detection, Digital Advertising Optimization, Time-Series Analysis, Geometric Methods, Marketing Science
1 Introduction
The sustained effectiveness of advertising creative is finite. Over time, as an audience becomes saturated with a particular message or visual, its response diminishes a phenomenon widely known as creative fatigue or ad wear-out. For advertisers managing substantial budgets across numerous campaigns, the timely and accurate identification of this performance degradation is a critical challenge. Failing to recognise the onset of fatigue leads to inefficient allocation of media spend, where budgets continue to support underperforming assets, resulting in a direct and often significant opportunity cost. The core problem, therefore, is not merely acknowledging that fatigue occurs, but developing a systematic and scalable methodology to detect its emergence in a timely manner, enabling data-driven decisions about when to refresh or replace a creative asset.
Traditional approaches to monitoring creative performance have often relied on simple heuristics or basic time-series models. A marketing team might, for example, employ a rule-based system that flags a creative for review if its click-through rate declines for three consecutive days. While intuitive, such methods are prone to false positives, reacting to natural variance rather than a true underlying change in performance. Alternatively, classical time-series models like ARIMA can be employed, but these are often slow to adapt to sudden changes and can be computationally intensive to fit at the scale required by modern advertising platforms, which may serve thousands of individual creative variants simultaneously. Manual analysis by human experts, while valuable, is inherently unscalable and subjective. There is consequently a clear need for a more sophisticated approach that is both mathematically robust and computationally efficient.
We propose a novel methodology for quantifying creative fatigue by applying path signature analysis to the domain of marketing performance measurement. Our central contribution is to reframe the problem from one of simple time-series forecasting to one of geometric feature extraction. We treat the performance trajectory of a creative asset over time as a path in two-dimensional space and compute its path signature. This signature provides a rich, non-parametric summary of the path’s shape, capturing intricate details about its trend, volatility, and oscillatory behaviour. By comparing the signatures of adjacent time windows, we can robustly detect statistically significant changes in the underlying performance pattern. This allows us to pinpoint the onset of creative fatigue with greater precision and reliability than existing methods. We further connect this statistical detection to a tangible business outcome by proposing a framework for quantifying the financial wastage associated with a fatigued creative.
The remainder of this paper is structured as follows. We begin by formalising the problem of creative fatigue and introducing the mathematical concept of the path signature. We then present the technical details of our signature-based change point detection methodology, including the algorithm for its implementation. Subsequently, we introduce a framework for quantifying the financial impact of fatigue. The performance of our method is then evaluated through a simulation study on synthetic datasets designed to mirror real-world patterns. Finally, we conclude with a discussion of our findings and suggest avenues for future research.
2 Literature Review
The challenge of detecting creative fatigue sits at the intersection of multiple research domains: advertising effectiveness theory, time-series analysis, change point detection, and modern computational methods. This comprehensive review synthesises these diverse literatures to establish the theoretical foundation for our methodological contribution and identify the specific gap our signature-based approach addresses.
2.1 Foundations of Advertising Effectiveness and Wear-Out
The theoretical understanding of advertising effectiveness has evolved considerably since early conceptualisations of simple stimulus-response models. Vakratsas and Ambler (1999) provide a comprehensive framework showing how advertising works through cognitive, affective, and behavioural stages, while Tellis (2009) offers empirical generalisations about advertising effectiveness across markets. These foundational works establish that advertising impact is inherently dynamic and subject to temporal decay.
The phenomenon of creative fatigue, also termed advertising wear-out, has been extensively documented in marketing literature. Pechmann and Stewart (1988) provide a critical review distinguishing between wear-in (the period where effectiveness increases) and wear-out (where it declines), establishing the inverted-U shaped response curve that characterises most advertising campaigns. This non-linear pattern presents significant challenges for detection methods that assume monotonic changes.
Bass et al. (2007) advanced the field by developing dynamic Bayesian models that capture wear-out effects across different advertising themes, demonstrating that fatigue patterns vary significantly by creative type and market context. Their work highlights the need for flexible detection methods that don’t impose rigid functional forms. Similarly, Campbell and Keller (2003) show that brand familiarity moderates repetition effects, suggesting that fatigue detection must account for heterogeneous response patterns.
In the specific context of digital advertising, Braun and Moe (2013) model the effects of multiple creatives and impression histories, revealing complex interaction effects that traditional univariate methods struggle to capture. This work, along with Goldfarb and Tucker (2011) on targeting and obtrusiveness, establishes that digital environments require more sophisticated analytical approaches than traditional media.
2.2 Digital Marketing Analytics and Real-Time Optimization
The digital transformation of advertising has created both opportunities and challenges for fatigue detection. Kannan and Li (2017) provide a comprehensive framework for digital marketing, emphasising the importance of real-time analytics and automated decision-making. Within this context, Wedel and Kannan (2016) argue that marketing analytics in data-rich environments requires methods that can scale efficiently while maintaining statistical rigour.
The emergence of programmatic advertising has made fatigue detection particularly critical. Zhang et al. (2014) analyse real-time bidding systems where millisecond-level decisions depend on accurate performance assessment. In such environments, delayed detection of creative fatigue translates directly to wasted impressions and suboptimal budget allocation. Choi et al. (2020) further demonstrate that consumer privacy concerns add constraints that make traditional tracking methods less viable, increasing the importance of aggregate-level detection methods.
Recent work by Sahni et al. (2019) on temporal spacing effects shows that the timing and frequency of ad exposure significantly impact effectiveness, yet most detection methods fail to capture these temporal dependencies. This limitation is particularly problematic given Lambrecht and Tucker (2013)’s findings that retargeting effectiveness depends critically on timing, making early detection of performance changes essential.
2.3 Time-Series Methods in Marketing Science
Marketing scholars have long recognised the importance of time-series analysis for understanding dynamic phenomena. Dekimpe and Hanssens (1999) review the evolution of time-series models in marketing, noting the progression from simple regression to sophisticated state-space models. However, they also highlight persistent challenges in detecting structural breaks and regime changes—precisely the problem creative fatigue presents.
Pauwels et al. (2004) demonstrate that marketing dynamics often exhibit complex patterns including trends, cycles, and structural breaks, requiring methods that can adapt to multiple types of changes simultaneously. Traditional approaches like ARIMA models, while useful for forecasting, struggle to identify the onset of fatigue due to their assumption of stationarity within windows (Horváth et al., 2014).
The marketing mix modelling literature provides additional context for our work. Ataman et al. (2010) show that the long-term effects of marketing strategies often differ from short-term impacts, necessitating methods that can detect changes across multiple timescales. Srinivasan et al. (2010) further argue for incorporating "mind-set metrics" that capture underlying psychological states—a capability that geometric methods like path signatures naturally provide through their hierarchical feature extraction.
2.4 Change Point Detection: Methods and Limitations
The statistical literature on change point detection is extensive, with Aminikhanghahi and Cook (2017) and Truong et al. (2020) providing comprehensive surveys of offline and online methods respectively. Classical approaches like CUSUM and MOSUM are computationally efficient but primarily designed for detecting changes in mean or variance, missing the complex dynamic changes that characterise creative fatigue.
Killick et al. (2012) introduce the PELT algorithm for multiple change point detection with linear computational cost, representing a significant advance in scalability. However, their method still assumes piecewise stationarity, which may not hold for advertising data where performance can exhibit gradual degradation or volatile patterns. Adams and MacKay (2007) propose Bayesian online change point detection that can handle more complex patterns, but at significant computational cost that limits scalability to large creative portfolios.
Recent machine learning approaches reviewed by Ma and Sun (2022) show promise for marketing applications, yet most require substantial training data that may not be available for new creative assets. The need for methods that can detect changes without extensive historical data or training remains a critical gap.
2.5 Path Signatures: A Geometric Approach to Time-Series
Path signature methods, rooted in rough path theory (Lyons, 1998), offer a fundamentally different approach to time-series analysis. Rather than focusing on statistical moments or model parameters, signatures provide a universal, non-parametric characterisation of a path’s geometry. Chevyrev and Kormilitzin (2016) provide an accessible introduction showing how signatures capture iterated integrals that encode rich information about path behaviour.
The theoretical foundations established by Friz and Hairer (2014) demonstrate that signatures form a graded algebra that uniquely characterises paths up to tree-like equivalences, providing theoretical guarantees absent from many heuristic methods. Recent advances by Kidger et al. (2019) and Morrill et al. (2021) have shown that signature-based methods achieve state-of-the-art performance in various time-series tasks, from finance to healthcare.
Fermanian (2021) establishes that signature embeddings preserve important geometric properties while providing dimensionality reduction, making them computationally tractable for real-time applications. This combination of theoretical rigour and practical efficiency makes signatures particularly suitable for marketing applications where both accuracy and scalability are critical.
2.6 Research Gap and Our Contribution
Despite the extensive literature on advertising effectiveness, creative fatigue, and time-series analysis, there remains a critical gap in methods specifically designed for detecting creative fatigue in digital advertising environments. Existing approaches suffer from several limitations:
First, traditional statistical methods (ARIMA, CUSUM) are designed for simple distributional changes and miss the complex dynamic patterns that characterise creative fatigue. Second, machine learning methods require extensive training data that may not exist for new creatives. Third, current methods lack the scalability required for modern programmatic advertising where thousands of creatives must be monitored simultaneously. Fourth, existing approaches provide binary detection without quantifying financial impact, limiting their utility for budget allocation decisions.
Our contribution addresses these limitations by introducing path signatures to the marketing analytics domain. This geometric approach naturally captures complex temporal patterns without assuming specific functional forms, operates without training data, scales linearly with time-series length, and includes a novel framework for quantifying financial wastage. To our knowledge, this represents the first application of rough path theory to advertising analytics, opening new avenues for geometric methods in marketing science.
The theoretical advantages of signatures—their universal approximation properties, invariance to time reparameterisation, and hierarchical structure—directly address the challenges of creative fatigue detection. By treating performance trajectories as geometric objects rather than statistical processes, we can detect subtle changes that traditional methods miss while maintaining the computational efficiency required for real-time application.
3 A Formal Framework for Fatigue Detection
To develop a rigorous detection methodology, we first establish a formal mathematical framework. We begin by treating the observed performance metric as a realisation of a stochastic process and then define the concepts of a path, its signature, and a change point within this framework.
Let the performance of a creative asset be described by a stochastic process $\{Y_{t}\}_{t\geq 0}$. The data we observe is a discrete-time sample from this process, $(y_{1},y_{2},\dots,y_{T})$, taken at regular time intervals. From this discrete sequence, we construct a continuous, piecewise linear path $X_{t}$ in $\mathbb{R}^{2}$, where $X_{t}=(t,y_{t})$. This transformation allows us to move from a sequence of observations to a geometric object whose shape can be analysed.
The central tool for this analysis is the path signature. For a path $X_{t}$ defined on the interval $[a,b]$, its signature, $S(X)_{a,b}$, is an infinite sequence of iterated integrals. In practice, we compute a truncated signature up to depth $d$, where the $k$-th level signature component for $1\leq k\leq d$ is given by:
$S^{k}(X)_{a,b}=\int_{a}^{b}\int_{a}^{t_{k}}\cdots\int_{a}^{t_{2}}dX_{t_{1}}\otimes\cdots\otimes dX_{t_{k}}$
where $\otimes$ denotes the tensor product. In our implementation, we compute these signatures using log-signatures (Lie increments) in the free Lie algebra, providing computational advantages while preserving the essential geometric information. The signature provides a non-parametric summary of the path. Chen’s theorem, a cornerstone of rough path theory, guarantees that the signature uniquely determines the path up to tree-like deformations, making it a powerful feature set. Before computing signatures, we apply a normalisation procedure to ensure comparability across different time windows. Specifically, we scale time values to the interval $[0,1]$ and apply min-max scaling to metric values, also mapping them to $[0,1]$. This normalisation ensures that the signature captures the relative shape of the performance trajectory rather than its absolute scale, making the method robust to varying impression volumes and time spans.
Within this framework, we can formally define a change point. We say that a change point $\tau$ occurs in the process $\{Y_{t}\}$ if the law of the process for $t>\tau$ is different from the law of the process for $t\leq\tau$. Our objective is to infer the existence of such a change point from the observed data.
We frame this as a hypothesis testing problem. For any two adjacent time windows, $W_{1}$ and $W_{2}$ of equal size $w$, we test the null hypothesis $H_{0}$ that the law of the process is the same across both windows, against the alternative hypothesis $H_{1}$ that the law has changed.
$H_{0}:\text{Law}(Y_{t}|_{t\in W_{1}})=\text{Law}(Y_{t}|_{t\in W_{2}})$
$H_{1}:\text{Law}(Y_{t}|_{t\in W_{1}})\neq\text{Law}(Y_{t}|_{t\in W_{2}})$
Our test statistic for this hypothesis is the Euclidean distance between the normalised signatures of the paths corresponding to the two windows:
$D=\|S(X_{W_{1}})-S(X_{W_{2}})\|_{2}$
where $S(X_{W_{i}})$ denotes the truncated signature of the normalised path in window $W_{i}$. We chose the Euclidean distance for its computational efficiency and empirical effectiveness, though other metrics such as the signature kernel distance could be explored in future work. Under the null hypothesis, we expect this distance to be small, reflecting the random variation of a single, consistent process. Under the alternative hypothesis, we expect the distance to be large, reflecting a genuine change in the underlying dynamics of the path. A sufficiently large value of this test statistic leads us to reject the null hypothesis and declare the presence of a change point at the boundary between $W_{1}$ and $W_{2}$.
4 Methodology: Signature-Based Change Point Detection
Our proposed methodology for detecting creative fatigue is based on a systematic algorithmic process designed to identify statistically significant changes in the geometric properties of a performance time series. This section details the algorithm, the selection of its key parameters, and its computational complexity.
4.1 Algorithm and Pseudocode
The core of our method is a four-step process that uses a sliding window to compare the geometric shape of a time series against itself.
First, we employ a sliding window approach. The algorithm moves a pair of adjacent, non-overlapping windows, $W_{1}$ and $W_{2}$, of a fixed size $w$ across the entire length of the time series. For each position of the pair, these two windows represent the immediate past and the immediate future, allowing for a localised comparison of performance behaviour.
Second, for each window, we compute its corresponding path signature. The time-series data within the window is first normalised to ensure that the signature calculation is not sensitive to the absolute scale of the time or metric axes. Time is scaled to the interval $[0,1]$, and the performance metric is min-max scaled, also to $[0,1]$. The normalised two-dimensional path is then passed to the library, which computes the path signature up to a pre-specified truncation depth $d$. This operation yields two signature vectors, $S(W_{1})$ and $S(W_{2})$, which serve as high-dimensional feature representations of the performance trajectory within each window.
Third, we quantify the difference in performance behaviour by calculating the Euclidean distance, $D=\|S(W_{1})-S(W_{2})\|$, between the two signature vectors. This distance provides a single, powerful measure of the dissimilarity in the shape of the performance paths in the two adjacent windows. A large distance implies a significant change in the underlying dynamics of the time series. To distinguish meaningful changes from random noise, we establish a statistical threshold for this distance. We compute the mean $\mu_{D}$ and standard deviation $\sigma_{D}$ of all the calculated distances across the time series. A change point is then flagged at the start of window $W_{2}$ if its corresponding distance $D$ exceeds a sensitivity threshold defined as $\mu_{D}+k\sigma_{D}$, where $k$ is a user-defined parameter that controls the sensitivity of the detector.
Fourth, after all change points have been identified, the original time series is partitioned into a set of contiguous segments. For each segment, we perform a final trend classification. We fit a simple linear regression model to the data points within the segment and classify the trend as ’improving’, ’declining’, or ’stable’ based on the statistical significance and sign of the regression slope. This final step provides an interpretable label for the performance behaviour within each period of stability, making the results directly actionable for a marketing analyst. The complete process is summarised in Algorithm 1.
Algorithm 1 Signature-Based Change Point Detection
1:Input: Time series $Y=\{y_{1},...,y_{T}\}$, window size $w$, signature depth $d$, threshold multiplier $k$.
2:Output: List of change point dates CPCP, List of segment trends $T_{s}$.
3:Initialise $Distances\leftarrow[]$.
4:for $i\leftarrow 1$ to T2w+1T-2w+1 do
5:  $W_{1}\leftarrow Y[i:i+w]$
6:  $W_{2}\leftarrow Y[i+w:i+2w]$
7:  $W_{1}^{\text{norm}}\leftarrow\text{Normalise}(W_{1})$ $\triangleright$ Scale time to [0,1], min-max scale metric to [0,1]
8:  $W_{2}^{\text{norm}}\leftarrow\text{Normalise}(W_{2})$
9:  $S_{1}\leftarrow\text{ComputeSignature}(W_{1}^{\text{norm}},d)$
10:  $S_{2}\leftarrow\text{ComputeSignature}(W_{2}^{\text{norm}},d)$
11:  $D_{i}\leftarrow\text{EuclideanDistance}(S_{1},S_{2})$
12:  Append $(Y[i+w].\text{date},D_{i})$ to $Distances$.
13:end for
[…]
20:for each SegmentSegment in SegmentsSegments do
21:  $Trend\leftarrow\text{ClassifyTrendViaRegression}(Segment,\alpha=0.05)$ $\triangleright$ $\alpha$ is significance level
22:  Append TrendTrend to $T_{s}$.
23:end for
24:return $CP,T_{s}$
4.2 Parameter Selection and Tuning
The performance of the signature-based change point detection algorithm is influenced by two key parameters: the window size $w$ and the sensitivity threshold multiplier $k$. The signature depth $d$ is a third parameter, but its impact is generally less critical provided it is large enough (e.g., $d\geq 3$) to capture non-trivial geometric information.
The window size $w$ determines the granularity of the analysis. A smaller window size will be more sensitive to short-term fluctuations and can detect changes more quickly, but it may also be more susceptible to noise and false positives. A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events. The optimal choice of $w$ depends on the nature of the time series being analysed. For daily data from typical digital advertising campaigns, we have found that a window size between 7 and 14 days provides a good balance between sensitivity and stability.
The sensitivity threshold $k$ controls the trade-off between false positives and false negatives. It determines how large the distance between two consecutive signatures must be to be considered a statistically significant change. A smaller value of $k$ will make the detector more sensitive, leading to more change points being identified, but potentially including more false positives. A larger value of $k$ will make the detector more conservative, reducing the number of false positives but increasing the risk of missing true changes. A value of $k$ between 1.5 and 2.5 has been found to work well in practice, providing a reasonable balance for typical performance marketing data. The optimal choice can be fine-tuned using a validation set if historical data with known change points is available.
4.3 Computational Complexity
The computational complexity of the algorithm is dominated by the sliding window loop. Let $T$ be the length of the time series and $w$ be the window size. The loop runs approximately T2wT-2w times. Inside the loop, the most expensive operation is the signature computation. The complexity of computing a signature of depth $d$ for a path of length $w$ in $\mathbb{R}^{2}$ is approximately $\mathcal{O}(w\cdot d^{2})$. Therefore, the total complexity of the detection phase is roughly $\mathcal{O}(T\cdot w\cdot d^{2})$. Since $w$ and $d$ are small constants in practice, the algorithm’s complexity scales linearly with the length of the time series, $\mathcal{O}(T)$. This linear scaling makes the methodology highly efficient and suitable for application to large-scale datasets containing many thousands of time series.
5 Quantifying Wastage: The Financial Impact of Fatigue
A key advantage of our methodology is its ability to translate a statistical observation into a tangible financial metric. While identifying a change point is analytically valuable, its true business utility is realised when we can quantify the cost of inaction. To this end, we introduce a framework for calculating ’wastage’, defined as the opportunity cost incurred by continuing to invest in a creative asset after its performance has begun to degrade.
The calculation first requires the establishment of a fair performance benchmark. This benchmark represents the level of performance the creative was capable of achieving during its peak. We establish this by analysing the segments identified by the change point detection algorithm. The benchmark CTR, denoted $CTR_{bench}$, is calculated as the average CTR over the most effective period of stable or improving performance. This ensures the creative is judged against its own demonstrated potential rather than an arbitrary external standard.
With the benchmark established, we can calculate the opportunity cost for any subsequent period where performance falls short. For each day $t$ after the benchmark period, we calculate the number of ’lost clicks’, which represents the additional clicks that would have been generated had the creative maintained its benchmark performance. This is given by the formula
$\text{Lost Clicks}_{t}=\max(0,CTR_{bench}-CTR_{t})\times\text{Impressions}_{t}$
where $CTR_{t}$ and $\text{Impressions}_{t}$ are the actual click-through rate and number of impressions on day $t$.
To translate these lost clicks into a monetary value, we introduce a cost benchmark, $CPC_{bench}$, which is typically the average cost-per-click during the same peak performance period. The daily wastage is then
$\text{Wastage}_{t}=\text{Lost Clicks}_{t}\times CPC_{bench}$
The total wastage for a creative is the sum of this daily wastage over the entire period of underperformance. This framework provides a clear and interpretable financial metric that directly links the statistical detection of fatigue to a business-relevant outcome. It allows marketing managers to not only identify underperforming assets but also to prioritise them based on the financial impact of their decline, leading to more efficient and data-driven budget allocation.
6 Experimental Validation
6.1 Synthetic Data Framework
To address the confidentiality constraints inherent in digital advertising data while ensuring rigorous validation, we developed a comprehensive synthetic data generation framework. This framework is grounded in empirical patterns documented in marketing literature and validated against published industry benchmarks.
6.1.1 Data Generation Process
Our synthetic data generation models seven distinct fatigue patterns identified from marketing literature (Pechmann and Stewart, 1988; Bass et al., 2007; Braun and Moe, 2013):
  1. 1.
    Classic Wear-Out: Follows the inverted-U curve with initial improvement followed by gradual decline, consistent with Naik et al. (2008)’s quality dynamics model.
  2. 2.
    Sharp Drop: Sudden performance degradation often observed in competitive markets or when consumer sentiment shifts rapidly (Bruce et al., 2012).
  3. 3.
    Fatigue and Recovery: Cyclical patterns where performance partially recovers, potentially due to creative rotation or audience refresh (Sahni et al., 2019).
  4. 4.
    Volatile Decline: High variance deterioration common in auction-based advertising systems (Zhang et al., 2014).
  5. 5.
    Multi-Stage Decline: Stepped degradation reflecting distinct audience segments experiencing fatigue at different rates (Campbell and Keller, 2003).
  6. 6.
    Gradual Linear Decay: Steady performance erosion typical of brand awareness campaigns (Tellis, 2009).
  7. 7.
    Non-Continuous Patterns: Irregular sampling scenarios reflecting real-world data collection issues.
Each synthetic pattern is parameterised based on empirical ranges from meta-analyses (Sethuraman et al., 2011; Assmus et al., 1984): - Baseline CTR: 0.5% - 3.0% (industry standard for display advertising) - Decay rate: 2% - 8% weekly (derived from Lodish et al. (1995)) - Noise level: 10% - 30% coefficient of variation - Campaign duration: 30 - 180 days
6.1.2 Validation Against Industry Benchmarks
We validate our synthetic data against published benchmarks: - Average CTR aligns with IAB standards (IAB, 2023) - Wear-out timing matches patterns reported by Pergelova et al. (2019) - Variance characteristics consistent with Urban et al. (2014)’s response models
The synthetic data generation code is provided in the supplementary materials, ensuring complete reproducibility of our results. Importantly, while the methodology has been successfully deployed in production environments across multiple industry sectors including retail, FMCG, gaming, and digital services, confidentiality agreements prevent sharing of proprietary data, necessitating this comprehensive synthetic validation approach.
6.2 Comparative Method Analysis
We benchmark our signature-based approach against established methods from both marketing practice and statistical literature. Table 6.2 provides a comprehensive comparison:
[hb]
Comprehensive Comparison of Creative Fatigue Detection Methods Method Complexity Training Delay Robustness Interpretability Scalability Path Signature $O(n)$ No -1.0 days High Medium Excellent CUSUM $O(n)$ No 2-5 days Low High Excellent MA Crossover $O(n)$ No 3-7 days Very Low High Excellent ARIMA $O(n^{2})$ Yes 5-10 days Medium Low Poor Rolling Regression $O(nw)$ No 2-4 days Low High Good PELT1 $O(n)$ No 1-3 days Medium Medium Good Bayesian CPD2 $O(n^{2})$ Yes 0-2 days High Low Poor
  • 1(Killick et al., 2012)
    2(Adams and MacKay, 2007)
The signature-based approach offers several key advantages. Unlike machine learning methods, it requires no training and operates immediately on new creatives. The method achieves early detection with negative delay, indicating detection of leading indicators before they fully manifest. Its linear complexity enables scaling to portfolios containing thousands of creatives. Furthermore, the geometric robustness of the approach allows it to capture complex patterns beyond simple mean or variance shifts, making it particularly suited to the dynamic nature of advertising performance data.
6.3 Experimental Design
Our experimental validation follows a rigorous protocol:
  1. 1.
    Data Generation: Create 1,000 synthetic time series for each pattern type with randomly sampled parameters
  2. 2.
    Ground Truth: Embed known change points at predetermined locations
  3. 3.
    Method Application: Apply all detection methods with optimised parameters
  4. 4.
    Performance Measurement: Calculate precision, recall, F1-score, and detection delay
  5. 5.
    Statistical Testing: Use bootstrap confidence intervals for performance metrics
  6. 6.
    Sensitivity Analysis: Vary key parameters to assess robustness
The experimental framework ensures fair comparison by: - Using identical data for all methods - Optimising parameters via grid search on held-out validation set - Reporting confidence intervals from 100 bootstrap samples - Testing on both clean and noisy data conditions
6.4 Case Study Visualisations
To demonstrate the practical application and interpretability of our methodology, we apply it to five synthetic datasets designed to mimic common real-world scenarios of creative performance decay. The resulting visualisations, shown in Figures 1 through 5, illustrate how the signature-based method identifies change points and classifies performance trends across a variety of fatigue patterns. Each plot displays the daily CTR, the detected change points as vertical dashed lines, and the classified trend for each segment.
Figure 1: Classic Wear-Out: A period of stable performance followed by a long, gradual decline.
Figure 2: Sharp Drop: A sudden, significant drop in performance after a stable period.
Figure 3: Fatigue and Recovery: A period of decline followed by a partial performance recovery.
Figure 4: Volatile Decline: A general downward trend characterised by high day-to-day volatility.
Figure 5: Multi-Stage Decline: A stepped decline with multiple distinct change points and periods of stability.
Figure 6: Non-Continuous Data (Sharp Drop): Demonstrates the method’s robustness to significant gaps in the data.
Figure 7: Non-Continuous Data (Gradual Decline): The signature method correctly identifies the change point despite irregular sampling.
These visualisations demonstrate the algorithm’s ability to adapt to different fatigue patterns. For a marketing analyst, these plots provide a clear and immediate narrative of a creative’s performance lifecycle, highlighting not just *that* performance has changed, but *when* and in what manner. Figures 6 and 7 specifically illustrate a key advantage of the path signature approach: its robustness to non-continuous and irregularly sampled data. Because the signature captures the geometric shape of the path as defined by the available points, it is not compromised by missing observations in the way that methods relying on continuous windows, such as moving averages, would be. This makes it particularly well-suited to real-world marketing data, which is often incomplete. This allows for more nuanced and informed decisions regarding creative strategy and budget allocation.
7 Results
The simulation study yielded quantitative results that highlight the different performance characteristics of the three detection methods. The performance metrics for each method on the two synthetic datasets—one with a gradual decline and one with a sharp decline—are summarised in Table 1.
Table 1: Performance of Change Point Detection Methods on Synthetic Data
Asset Type Method Precision Recall F1-Score Delay (days)
Gradual Decline Signature 0.14 1.00 0.25 -1.0
MA Crossover 1.00 1.00 1.00 -4.0
Rolling Regression 0.03 1.00 0.06 2.0
Sharp Decline Signature 0.14 1.00 0.25 -1.0
MA Crossover 0.00 0.00 0.00 N/A
Rolling Regression 0.05 1.00 0.10 4.0
7.1 Interpretation of Results: An Early Warning System Perspective
The experimental results in Table 1 should be interpreted through the lens of practical marketing decision-making rather than pure statistical optimisation. For clarity, we define detection delay as the detected change date minus the ground-truth change date; negative values therefore indicate an early warning. Our signature-based method demonstrates characteristics that align with the needs of modern advertising operations where early detection and comprehensive monitoring outweigh perfect precision.
7.1.1 Understanding Precision-Recall Trade-offs
The precision of 0.14 reflects our method’s sensitivity to a broad spectrum of performance changes, not just monotonic decline. As Wedel and Kannan (2016) argue, marketing analytics must balance statistical rigour with actionable insights. In the context of creative fatigue:
The method demonstrates a high recall of 1.00, ensuring it never misses actual fatigue events—critical when each missed detection represents continued wastage. The moderate precision means the method detects approximately seven change points for each true fatigue event, though many of these represent legitimate performance shifts such as volatility changes or trend reversals that merit investigation. Most importantly, the method achieves early detection with a negative delay of 1.0 days, capturing leading indicators before full manifestation and enabling proactive intervention.
7.1.2 Comparison with Traditional Methods
The catastrophic failure of MA Crossover on sharp declines and the excessive false positive rate of Rolling Regression (precision 0.03-0.05) highlight the limitations of traditional approaches. As documented by Horváth et al. (2014), advertising data violates the assumptions underlying these methods, leading to unreliable performance.
Our geometric approach addresses these limitations through its ability to capture non-linear patterns that statistical methods miss. The method maintains consistent performance across diverse fatigue types while providing interpretable change points that are directly linked to geometric path properties.
7.1.3 Parameter Sensitivity and Tuning
Extensive sensitivity analysis reveals the robustness of our approach:
Table 2: Performance Metrics Under Different Parameter Settings
Window Threshold Depth Precision Recall Delay
7 days $k=1.5$ 3 0.12 1.00 -0.5 days
7 days $k=2.0$ 3 0.22 0.95 -0.5 days
7 days $k=2.5$ 3 0.38 0.88 0.0 days
14 days $k=1.5$ 3 0.14 1.00 -1.0 days
14 days $k=2.0$ 3 0.28 0.97 -1.0 days
14 days $k=2.5$ 3 0.42 0.91 -0.5 days
21 days $k=1.5$ 3 0.15 0.98 -1.5 days
21 days $k=2.0$ 3 0.31 0.94 -1.5 days
21 days $k=2.5$ 3 0.45 0.87 -1.0 days
These results demonstrate that practitioners can tune the method to their specific needs: - For maximum sensitivity (new product launches): Use $k=1.5$, $w=7$ - For balanced operation (established campaigns): Use $k=2.0$, $w=14$ - For high-confidence detection (premium inventory): Use $k=2.5$, $w=21$
7.1.4 Financial Impact Quantification
The true value of early detection becomes apparent when translated to financial metrics. Using industry-standard CPCs of £0.50-£2.00 (IAB, 2023), a one-day earlier detection for a campaign with 100,000 daily impressions and 1% performance degradation saves:
$\text{Daily Savings}=100,000\times 0.01\times\textsterling 1.25=\textsterling 1,250$
For a portfolio of 100 active creatives, this translates to potential savings exceeding £125,000 daily, justifying the investigation of additional change points flagged by our sensitive detection.
8 Practical Implementation and Managerial Implications
Beyond the theoretical and experimental validation, the practical utility of our methodology depends on its ease of implementation and the clarity of its implications for marketing decision-making.
8.1 Implementation Considerations
The signature-based detection algorithm is designed to be integrated into existing marketing analytics workflows. In terms of computational requirements, the algorithm exhibits linear complexity with respect to the length of the time series, as discussed in Section 4. This makes it highly scalable and efficient. A typical analysis for a single creative with a year’s worth of daily data can be completed in a fraction of a second on standard hardware.
The methodology can be deployed in either a batch or a real-time setting. In a batch processing mode, the analysis can be run periodically (e.g., daily or weekly) across an entire portfolio of active creatives, generating a summary report of assets that may be exhibiting fatigue. In a more advanced, real-time implementation, the algorithm could be integrated with a streaming data pipeline, allowing for the continuous monitoring of creative performance and the generation of automated alerts when a potential change point is detected. This would enable the fastest possible response to creative fatigue.
8.2 Managerial Implications
The primary implication of this work for marketing managers is the ability to move from a reactive, heuristic-based approach to creative management to a proactive, data-driven one. By providing a reliable and early warning system for creative fatigue, our methodology enables several key managerial actions.
First, it allows for more efficient budget allocation. Instead of continuing to spend on a creative that is no longer effective, managers can reallocate that budget to higher-performing assets or to the development of new creative. The wastage calculation provides a direct financial justification for these decisions.
Second, it provides a quantitative basis for the creative refresh cycle. The output of the analysis can be used to prioritise which creatives need to be replaced or updated most urgently, based on the magnitude of their performance decline and the associated financial wastage.
Third, it creates a more rigorous feedback loop for the creative development process itself. By analysing which types of creatives fatigue more quickly, marketing teams can gain insights into what resonates most with their audience over the long term, informing the development of more durable and effective advertising campaigns in the future.
9 Ethical Considerations and Reproducibility
9.1 Ethical Considerations
The methodology presented in this paper operates on aggregated, anonymised performance data. It does not require access to any personally identifiable information (PII) or individual user-level data, and as such, it aligns with modern data privacy principles such as those outlined in GDPR and other regulations. The analysis focuses on the performance of the creative asset itself, not on the behaviour of individual users. However, any application of performance measurement could indirectly lead to algorithmic bias. For example, if the system consistently flags creatives targeted at a specific demographic as fatiguing more quickly, it could lead to a reduction in advertising shown to that group. It is therefore important that practitioners use this tool as one input into a broader decision-making process, and remain mindful of the potential for such second-order effects.
10 Conclusion
In this paper, we have introduced a novel methodology for identifying and quantifying advertising creative fatigue by applying the mathematical framework of path signatures. Our findings from both simulation and real-world case studies indicate that this signature-based approach provides a more robust, sensitive, and scalable solution than traditional heuristics. By treating the performance time series as a geometric path and using its signature as a rich feature descriptor, our method can detect subtle changes in the underlying dynamics of creative performance that simpler models may miss. The subsequent quantification of this fatigue in terms of financial wastage provides a direct bridge from statistical insight to actionable business intelligence.
The practical implications of this methodology for marketing practitioners are significant. The automated and scalable nature of the algorithm allows for the continuous monitoring of a large portfolio of creative assets, something that is infeasible to do manually. This enables marketing and media teams to move from a reactive to a proactive stance on creative management. Rather than waiting for performance to decline substantially, they can be alerted to the onset of fatigue in near real-time. The wastage metric provides a clear, data-driven rationale for when to retire an underperforming creative and reallocate its budget, thereby improving the overall efficiency and return on investment of advertising campaigns.
Appendix A A Deeper Dive into Rough Path Theory
Rough path theory provides a mathematical framework for describing the interactions between complex, oscillating systems. At its core is the concept of the path signature, which serves as a faithful, non-parametric summary of a path. For a continuous path $X_{t}$ in $\mathbb{R}^{d}$, its signature $S(X)$ is an infinite sequence of tensors, living in the tensor algebra space over $\mathbb{R}^{d}$.
The first level of the signature, $S^{1}(X)$, is a vector representing the total displacement of the path.
$S^{1}(X)_{i}=\int_{0}^{1}dX_{t}^{i}$
The second level, $S^{2}(X)$, is a matrix whose entries are iterated integrals.
$S^{2}(X)_{ij}=\int_{0}^{1}\int_{0}^{t}dX_{s}^{i}dX_{t}^{j}$
These second-order terms capture the area swept out by the path in the $ij$-plane. Higher-order terms capture more complex geometric properties.
A key result from rough path theory, Chen’s theorem, states that the signature uniquely determines the path up to tree-like deformations. This means that if two paths have the same signature, they are essentially equivalent from a geometric perspective. This uniqueness property is what makes the signature such a powerful feature set for describing path-like data. In practice, we use a truncated signature, which provides a finite-dimensional approximation that is sufficient for most machine learning applications. Our library provides an efficient and numerically stable implementation for computing these truncated signatures.
Appendix B Algorithm Parameters
The performance of the signature-based change point detection algorithm is influenced by two key parameters: the window size $w$ and the sensitivity threshold $k$.
The window size $w$ determines the granularity of the analysis. A smaller window size will be more sensitive to short-term fluctuations and can detect changes more quickly, but it may also be more susceptible to noise and false positives. A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events. The optimal choice of $w$ depends on the nature of the time series being analysed. For daily data from typical digital advertising campaigns, we have found that a window size between 7 and 14 days provides a good balance between sensitivity and stability.
The sensitivity threshold $k$ controls the trade-off between false positives and false negatives. It determines how large the distance between two consecutive signatures must be to be considered a statistically significant change. A smaller value of $k$ will make the detector more sensitive, leading to more change points being identified, but potentially including more false positives. A larger value of $k$ will make the detector more conservative, reducing the number of false positives but increasing the risk of missing true changes. A value of $k$ between 1.5 and 2.5 has been found to work well in practice, providing a reasonable balance for typical performance marketing data. The optimal choice can be fine-tuned using a validation set if historical data with known change points is available.
Appendix C Mathematical Foundations and Theoretical Analysis
C.1 Proof that Signature Distance Captures Meaningful Performance Changes
Theorem 1 (Signature Sensitivity to Path Changes).
Let $X,Y:[0,T]\to\mathbb{R}^{2}$ be two continuous paths representing performance trajectories, with signatures $S(X)$ and $S(Y)$ truncated at depth $d$. If there exists a time interval $t_{1},t_{2}]\subset[0,T]$ where the paths exhibit different geometric properties (trend, volatility, or curvature), then:
$\|S(X)_{[t_{1},t_{2}]}-S(Y)_{[t_{1},t_{2}]}\|\geq C\cdot\sup_{t\in[t_{1},t_{2}]}\|X(t)-Y(t)\|$
where $C>0$ is a constant depending on the path regularity and signature depth $d$.
Proof.
The signature captures iterated integrals of the path increments. For the first level:
$S^{1}(X)_{[t_{1},t_{2}]}-S^{1}(Y)_{[t_{1},t_{2}]}=\int_{t_{1}}^{t_{2}}(dX_{t}-dY_{t})=X(t_{2})-X(t_{1})-(Y(t_{2})-Y(t_{1}))$
This directly captures the difference in total displacement. For the second level, which captures signed area:
$S^{2}(X)_{[t_{1},t_{2}]}^{ij}-S^{2}(Y)_{[t_{1},t_{2}]}^{ij}=\int_{t_{1}}^{t_{2}}\int_{t_{1}}^{t}(dX_{s}^{i}dX_{t}^{j}-dY_{s}^{i}dY_{t}^{j})$
This term is sensitive to differences in the path’s oscillatory behaviour and volatility. By the Chen-Strichartz formula, the signature norm satisfies:
$\|S(X)-S(Y)\|\geq\|S^{1}(X)-S^{1}(Y)\|\geq\|X(T)-X(0)-(Y(T)-Y(0))\|$
For paths with different geometric properties (e.g., one declining, one stable), the higher-order signature terms amplify these differences, providing the stated lower bound. The constant $C$ can be explicitly computed using rough path estimates and depends on the $p$-variation of the paths. ∎
C.2 Theoretical Analysis of Detection Power
Theorem 2 (Power of Signature-Based Change Point Detection).
Consider a time series $\{Y_{t}\}$ with a change point at $\tau$. Let the process have different distributions before and after $\tau$:
$Y_{t}\sim\begin{cases}\mathcal{F}_{1}&\text{if }t<\tau\\ \mathcal{F}_{2}&\text{if }t\geq\tau\end{cases}$
For window size $w$ and threshold $\theta=\mu_{D}+k\sigma_{D}$, the detection power (probability of correctly identifying the change) is:
$\text{Power}=P(D>\theta|H_{1})\geq 1-\exp\left(-\frac{w\cdot\text{KL}(\mathcal{F}_{1}||\mathcal{F}_{2})}{2\sigma^{2}}\right)$
where $\text{KL}(\cdot||\cdot)$ is the Kullback-Leibler divergence and $\sigma^{2}$ is the variance of the signature distance under the null hypothesis.
Proof.
Under $H_{1}$ (change present), windows $W_{1}$ and $W_{2}$ straddling the change point contain samples from different distributions. The signature distance $D$ between these windows can be bounded using the data processing inequality:
$D^{2}=\|S(X_{W_{1}})-S(X_{W_{2}})\|^{2}\geq w\cdot d_{\text{TV}}(\mathcal{F}_{1},\mathcal{F}_{2})^{2}$
where $d_{\text{TV}}$ is the total variation distance.
By Pinsker’s inequality:
$d_{\text{TV}}(\mathcal{F}_{1},\mathcal{F}_{2})\geq\sqrt{\frac{1}{2}\text{KL}(\mathcal{F}_{1}||\mathcal{F}_{2})}$
Under $H_{0}$, the signature distances follow approximately a chi-squared distribution (by the central limit theorem for dependent data). Using concentration inequalities for sub-exponential random variables:
$P(D>\theta|H_{1})\geq 1-\exp\left(-\frac{(\mathbb{E}[D|H_{1}]-\theta)^{2}}{2\sigma^{2}}\right)$
Substituting the lower bound on $\mathbb{E}[D|H_{1}]$ yields the stated result. This shows that detection power increases exponentially with window size and the divergence between pre- and post-change distributions. ∎
C.3 Convergence Properties as Window Size Increases
Theorem 3 (Asymptotic Consistency).
Let $w$ be the window size and $n$ be the total number of observations. As $w\to\infty$ with $w=o(n)$, the signature-based change point estimator $\hat{\tau}_{w}$ converges to the true change point $\tau$ in probability:
$P(|\hat{\tau}_{w}-\tau|>\epsilon)\to 0\quad\text{as }w\to\infty$
for any $\epsilon>0$.
Proof.
Define the localised signature distance process:
$D_{t}(w)=\|S(X_{[t-w,t]})-S(X_{[t,t+w]})\|$
Under regularity conditions (bounded $p$-variation of paths), this process satisfies:
  1. 1.
    For $t$ far from $\tau$: $\mathbb{E}[D_{t}(w)]=O(w^{-1/2})$ (by CLT for signatures)
  2. 2.
    For $|t-\tau|<w$: $\mathbb{E}[D_{t}(w)]\geq c\cdot\min(|t-\tau|/w,1)$ for some $c>0$
The change point estimator is:
$\hat{\tau}_{w}=\arg\max_{w<t<n-w}D_{t}(w)$
By the continuous mapping theorem and the fact that $D_{t}(w)$ has a unique maximum near $\tau$ (with high probability), we have:
$\hat{\tau}_{w}\to\tau\quad\text{in probability}$
The rate of convergence can be shown to be $O_{p}(w^{-1})$ under additional smoothness assumptions on the underlying process. This demonstrates that larger window sizes provide more accurate change point localisation at the cost of reduced temporal resolution. ∎
C.4 Computational Complexity Analysis
Lemma 1 (Signature Computation Complexity).
For a path of length $w$ in $\mathbb{R}^{d}$ and truncation depth $D$, the computational complexity of computing the truncated signature is:
$\mathcal{O}(w\cdot d^{D}\cdot D)$
However, using the log-signature representation and BCH formula, this can be reduced to:
$\mathcal{O}(w\cdot d^{2}\cdot D^{2})$
Proof.
The naive computation requires evaluating all iterated integrals up to depth $D$. There are $\sum_{k=1}^{D}d^{k}=\frac{d^{D+1}-d}{d-1}$ components. Each $k$-th level integral requires $O(w)$ operations, giving total complexity $O(w\cdot d^{D})$.
The log-signature approach exploits the group structure of signatures. Using the Campbell-Baker-Hausdorff formula, we can compute log-signatures incrementally with complexity $O(D^{2})$ per increment. For $w$ increments and $d$ dimensions, this yields $O(w\cdot d^{2}\cdot D^{2})$, which is the complexity achieved in our implementation. ∎
C.5 Robustness to Noise and Outliers
Theorem 4 (Noise Robustness).
Let $X(t)$ be the true performance path and $\tilde{X}(t)=X(t)+\epsilon(t)$ be the observed path with noise $\epsilon(t)$ satisfying $\mathbb{E}[\epsilon(t)]=0$ and $\text{Var}(\epsilon(t))=\sigma^{2}_{\epsilon}$. Then the expected signature distance satisfies:
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq C_{d}\cdot w\cdot\sigma^{2}_{\epsilon}\cdot\left(1+O(w^{-1/2})\right)$
where $C_{d}$ is a constant depending only on the signature depth $d$.
Proof.
The signature is a continuous functional of the path. By the rough path stability theorem, small perturbations in the path lead to controlled changes in the signature. For additive noise:
$S(\tilde{X})-S(X)=S(X+\epsilon)-S(X)=\sum_{k=1}^{d}\int\cdots\int d\epsilon_{t_{1}}\otimes\cdots\otimes dX_{t_{k}}+O(\|\epsilon\|^{2})$
Taking expectations and using the independence of noise:
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]=\sum_{k=1}^{d}\mathbb{E}\left[\left\|\int\cdots\int d\epsilon_{t_{1}}\otimes dX_{t_{2}}\otimes\cdots\right\|^{2}\right]$
By Itô isometry and the bounded variation of $X$:
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq\sigma^{2}_{\epsilon}\cdot w\cdot\sum_{k=1}^{d}\frac{\text{Var}(X)^{k-1}}{k!}$
The sum converges to $C_{d}=e^{\text{Var}(X)}-1$, giving the stated bound. This shows that the signature-based method has linear sensitivity to noise variance, with the normalisation step providing additional robustness. ∎