### v3

# A Path Signature Framework for Detecting Creative Fatigue in Digital Advertising

Abstract
This paper introduces a signature-based framework for detecting advertising creative fatigue using path signatures, a geometric representation from rough path theory. Creative fatigue—the degradation of creative effectiveness under repeated exposure—is operationally important in digital marketing because delayed detection can translate directly into avoidable opportunity cost. We reframe fatigue monitoring as a geometric change detection problem: advertising performance trajectories are embedded as paths and represented by truncated (log-)signatures, enabling detection of changes in trend, volatility, and non-linear dynamics beyond simple mean or variance shifts. We further connect statistical detection to managerial decision-making via an explicit quantification of performance loss relative to a benchmark period.
Because proprietary production data cannot be released, we evaluate the proposed framework on a synthetic panel dataset designed to mimic realistic impression volumes and noisy day-to-day CTR dynamics. We define observed CTR as the realised binomial rate $CTR_{t}:=C_{t}/I_{t}$ using daily clicks $C_{t}$ and impressions $I_{t}$. The accompanying CSV also contains a pre-computed CTR field (e.g., due to rounding or upstream derivation), but all modelling and evaluation in this paper use $C_{t}/I_{t}$. Crucially, the dataset does not include injected changepoints; we therefore define an operational ground truth for “fatigue onset” based on a noise-robust CTR estimate and a sustained deterioration relative to a recent-best baseline. We report lead-time (early warning) and alert-burden metrics under this operational definition, and provide a sensitivity analysis over the detector’s primary tuning parameters. The methodology scales linearly in time-series length for fixed signature depth and is suitable for monitoring large creative portfolios.
Keywords: Marketing Analytics, Creative Fatigue Detection, Path Signatures, Rough Path Theory, Change Point Detection, Digital Advertising Optimization, Time-Series Analysis, Geometric Methods, Marketing Science
1 Introduction
The sustained effectiveness of advertising creative is finite. Over time, as an audience becomes saturated with a particular message or visual, its response diminishes a phenomenon widely known as creative fatigue or ad wear-out. For advertisers managing substantial budgets across numerous campaigns, the timely and accurate identification of this performance degradation is a critical challenge. Failing to recognise the onset of fatigue leads to inefficient allocation of media spend, where budgets continue to support underperforming assets, resulting in a direct and often significant opportunity cost. The core problem, therefore, is not merely acknowledging that fatigue occurs, but developing a systematic and scalable methodology to detect its emergence in a timely manner, enabling data-driven decisions about when to refresh or replace a creative asset.
Traditional approaches to monitoring creative performance have often relied on simple heuristics or basic time-series models. A marketing team might, for example, employ a rule-based system that flags a creative for review if its click-through rate declines for three consecutive days. While intuitive, such methods are prone to false positives, reacting to natural variance rather than a true underlying change in performance. Alternatively, classical time-series models like ARIMA can be employed, but these are often slow to adapt to sudden changes and can be computationally intensive to fit at the scale required by modern advertising platforms, which may serve thousands of individual creative variants simultaneously. Manual analysis by human experts, while valuable, is inherently unscalable and subjective. There is consequently a clear need for a more sophisticated approach that is both mathematically robust and computationally efficient.
We propose a novel methodology for quantifying creative fatigue by applying path signature analysis to the domain of marketing performance measurement. Our central contribution is to reframe the problem from one of simple time-series forecasting to one of geometric feature extraction. We treat the performance trajectory of a creative asset over time as a path in two-dimensional space and compute its path signature. This signature provides a rich, non-parametric summary of the path’s shape, capturing intricate details about its trend, volatility, and oscillatory behaviour. By comparing the signatures of adjacent time windows, we can robustly detect statistically significant changes in the underlying performance pattern. This allows us to pinpoint the onset of creative fatigue with greater precision and reliability than existing methods. We further connect this statistical detection to a tangible business outcome by proposing a framework for quantifying the financial wastage associated with a fatigued creative.
Our contributions are fourfold. First, we introduce a signature distance statistic for monitoring creative performance time series, with an explicit algorithm and complexity analysis suitable for large-scale deployment. Second, we formalise a practical mapping from detected regime changes to an interpretable fatigue signal via segment-level trend classification. Third, we propose a benchmark-relative quantification of performance loss that links detection outputs to managerial prioritisation. Fourth, we provide a technically transparent evaluation protocol for noisy daily CTR data by defining an operational ground truth for fatigue onset and reporting lead-time and alert-burden metrics alongside sensitivity analyses.
The remainder of this paper is structured as follows. We begin by formalising the problem of creative fatigue and introducing the mathematical concept of the path signature. We then present the technical details of our signature-based change point detection methodology, including the algorithm for its implementation. Subsequently, we introduce a framework for quantifying the financial impact of fatigue. The performance of our method is then evaluated through a simulation study on synthetic datasets designed to mirror real-world patterns. Finally, we conclude with a discussion of our findings and suggest avenues for future research.
2 Literature Review
The challenge of detecting creative fatigue sits at the intersection of multiple research domains: advertising effectiveness theory, time-series analysis, change point detection, and modern computational methods. This comprehensive review synthesises these diverse literatures to establish the theoretical foundation for our methodological contribution and identify the specific gap our signature-based approach addresses.
2.1 Foundations of Advertising Effectiveness and Wear-Out
The theoretical understanding of advertising effectiveness has evolved considerably since early conceptualisations of simple stimulus-response models. Vakratsas and Ambler (1999) provide a comprehensive framework showing how advertising works through cognitive, affective, and behavioural stages, while Tellis (2009) offers empirical generalisations about advertising effectiveness across markets. These foundational works establish that advertising impact is inherently dynamic and subject to temporal decay.
The phenomenon of creative fatigue, also termed advertising wear-out, has been extensively documented in marketing literature. Pechmann and Stewart (1988) provide a critical review distinguishing between wear-in (the period where effectiveness increases) and wear-out (where it declines), establishing the inverted-U shaped response curve that characterises most advertising campaigns. This non-linear pattern presents significant challenges for detection methods that assume monotonic changes.
Bass et al. (2007) advanced the field by developing dynamic Bayesian models that capture wear-out effects across different advertising themes, demonstrating that fatigue patterns vary significantly by creative type and market context. Their work highlights the need for flexible detection methods that don’t impose rigid functional forms. Similarly, Campbell and Keller (2003) show that brand familiarity moderates repetition effects, suggesting that fatigue detection must account for heterogeneous response patterns.
In the specific context of digital advertising, Braun and Moe (2013) model the effects of multiple creatives and impression histories, revealing complex interaction effects that traditional univariate methods struggle to capture. This work, along with Goldfarb and Tucker (2011) on targeting and obtrusiveness, establishes that digital environments require more sophisticated analytical approaches than traditional media.
2.2 Digital Marketing Analytics and Real-Time Optimization
The digital transformation of advertising has created both opportunities and challenges for fatigue detection. Kannan and Li (2017) provide a comprehensive framework for digital marketing, emphasising the importance of real-time analytics and automated decision-making. Within this context, Wedel and Kannan (2016) argue that marketing analytics in data-rich environments requires methods that can scale efficiently while maintaining statistical rigour.
The emergence of programmatic advertising has made fatigue detection particularly critical. Zhang et al. (2014) analyse real-time bidding systems where millisecond-level decisions depend on accurate performance assessment. In such environments, delayed detection of creative fatigue translates directly to wasted impressions and suboptimal budget allocation. Choi et al. (2020) further demonstrate that consumer privacy concerns add constraints that make traditional tracking methods less viable, increasing the importance of aggregate-level detection methods.
Recent work by Sahni et al. (2019) on temporal spacing effects shows that the timing and frequency of ad exposure significantly impact effectiveness, yet most detection methods fail to capture these temporal dependencies. This limitation is particularly problematic given Lambrecht and Tucker (2013)’s findings that retargeting effectiveness depends critically on timing, making early detection of performance changes essential.
2.3 Time-Series Methods in Marketing Science
Marketing scholars have long recognised the importance of time-series analysis for understanding dynamic phenomena. Dekimpe and Hanssens (1999) review the evolution of time-series models in marketing, noting the progression from simple regression to sophisticated state-space models. However, they also highlight persistent challenges in detecting structural breaks and regime changes—precisely the problem creative fatigue presents.
Pauwels et al. (2004) demonstrate that marketing dynamics often exhibit complex patterns including trends, cycles, and structural breaks, requiring methods that can adapt to multiple types of changes simultaneously. Traditional approaches like ARIMA models, while useful for forecasting, struggle to identify the onset of fatigue due to their assumption of stationarity within windows (Horváth et al., 2014).
The marketing mix modelling literature provides additional context for our work. Ataman et al. (2010) show that the long-term effects of marketing strategies often differ from short-term impacts, necessitating methods that can detect changes across multiple timescales. Srinivasan et al. (2010) further argue for incorporating "mind-set metrics" that capture underlying psychological states—a capability that geometric methods like path signatures naturally provide through their hierarchical feature extraction.
2.4 Change Point Detection: Methods and Limitations
The statistical literature on change point detection is extensive, with Aminikhanghahi and Cook (2017) and Truong et al. (2020) providing comprehensive surveys of offline and online methods respectively. Classical approaches like CUSUM and MOSUM are computationally efficient but primarily designed for detecting changes in mean or variance, missing the complex dynamic changes that characterise creative fatigue.
Killick et al. (2012) introduce the PELT algorithm for multiple change point detection with linear computational cost, representing a significant advance in scalability. However, their method still assumes piecewise stationarity, which may not hold for advertising data where performance can exhibit gradual degradation or volatile patterns. Adams and MacKay (2007) propose Bayesian online change point detection that can handle more complex patterns, but at significant computational cost that limits scalability to large creative portfolios.
Recent machine learning approaches reviewed by Ma and Sun (2022) show promise for marketing applications, yet most require substantial training data that may not be available for new creative assets. The need for methods that can detect changes without extensive historical data or training remains a critical gap.
2.5 Path Signatures: A Geometric Approach to Time-Series
Path signature methods, rooted in rough path theory (Lyons, 1998), offer a fundamentally different approach to time-series analysis. Rather than focusing on statistical moments or model parameters, signatures provide a universal, non-parametric characterisation of a path’s geometry. Chevyrev and Kormilitzin (2016) provide an accessible introduction showing how signatures capture iterated integrals that encode rich information about path behaviour.
The theoretical foundations established by Friz and Hairer (2014) demonstrate that signatures form a graded algebra that uniquely characterises paths up to tree-like equivalences, providing theoretical guarantees absent from many heuristic methods. Recent advances by Kidger et al. (2019) and Morrill et al. (2021) have shown that signature-based methods achieve state-of-the-art performance in various time-series tasks, from finance to healthcare.
Fermanian (2021) establishes that signature embeddings preserve important geometric properties while providing dimensionality reduction, making them computationally tractable for real-time applications. This combination of theoretical rigour and practical efficiency makes signatures particularly suitable for marketing applications where both accuracy and scalability are critical.
2.6 Research Gap and Our Contribution
Despite the extensive literature on advertising effectiveness, creative fatigue, and time-series analysis, there remains a critical gap in methods specifically designed for detecting creative fatigue in digital advertising environments. Existing approaches suffer from several limitations:
First, traditional statistical methods (ARIMA, CUSUM) are designed for simple distributional changes and miss the complex dynamic patterns that characterise creative fatigue. Second, machine learning methods require extensive training data that may not exist for new creatives. Third, current methods lack the scalability required for modern programmatic advertising where thousands of creatives must be monitored simultaneously. Fourth, existing approaches provide binary detection without quantifying financial impact, limiting their utility for budget allocation decisions.
Our contribution addresses these limitations by introducing path signatures to the marketing analytics domain. This geometric approach naturally captures complex temporal patterns without assuming specific functional forms, operates without training data, scales linearly with time-series length, and includes a novel framework for quantifying financial wastage. To our knowledge, this represents the first application of rough path theory to advertising analytics, opening new avenues for geometric methods in marketing science.
The theoretical advantages of signatures—their universal approximation properties, invariance to time reparameterisation, and hierarchical structure—directly address the challenges of creative fatigue detection. By treating performance trajectories as geometric objects rather than statistical processes, we can detect subtle changes that traditional methods miss while maintaining the computational efficiency required for real-time application.
3 A Formal Framework for Fatigue Detection
To develop a rigorous detection methodology, we first establish a formal mathematical framework. We begin by treating the observed performance metric as a realisation of a stochastic process and then define the concepts of a path, its signature, and a change point within this framework.
Let the performance of a creative asset be described by a stochastic process $\{Y_{t}\}_{t\geq 0}$. The data we observe is a discrete-time sample from this process, $(y_{1},y_{2},\dots,y_{T})$, taken at regular time intervals. From this discrete sequence, we construct a continuous, piecewise linear path $X_{t}$ in $\mathbb{R}^{2}$, where $X_{t}=(t,y_{t})$. This transformation allows us to move from a sequence of observations to a geometric object whose shape can be analysed.
The central tool for this analysis is the path signature. For a path $X_{t}$ defined on the interval $[a,b]$, its signature, $S(X)_{a,b}$, is an infinite sequence of iterated integrals. In practice, we compute a truncated signature up to depth $d$, where the $k$-th level signature component for $1\leq k\leq d$ is given by:
$S^{k}(X)_{a,b}=\int_{a}^{b}\int_{a}^{t_{k}}\cdots\int_{a}^{t_{2}}dX_{t_{1}}\otimes\cdots\otimes dX_{t_{k}}$
where $\otimes$ denotes the tensor product. In our implementation, we compute these signatures using log-signatures (Lie increments) in the free Lie algebra, providing computational advantages while preserving the essential geometric information. The signature provides a non-parametric summary of the path. Chen’s theorem, a cornerstone of rough path theory, guarantees that the signature uniquely determines the path up to tree-like deformations, making it a powerful feature set. Before computing signatures, we apply a normalisation procedure to ensure comparability across different time windows. Specifically, we scale time values to the interval $[0,1]$ and apply min-max scaling to metric values, also mapping them to $[0,1]$. This normalisation ensures that the signature captures the relative shape of the performance trajectory rather than its absolute scale, making the method robust to varying impression volumes and time spans.
Within this framework, we can formally define a change point. We say that a change point $\tau$ occurs in the process $\{Y_{t}\}$ if the law of the process for $t>\tau$ is different from the law of the process for $t\leq\tau$. Our objective is to infer the existence of such a change point from the observed data.
We frame this as a hypothesis testing problem. For any two adjacent time windows, $W_{1}$ and $W_{2}$ of equal size $w_{d}$, we test the null hypothesis $H_{0}$ that the law of the process is the same across both windows, against the alternative hypothesis $H_{1}$ that the law has changed.
$H_{0}:\text{Law}(Y_{t}|_{t\in W_{1}})=\text{Law}(Y_{t}|_{t\in W_{2}})$
$H_{1}:\text{Law}(Y_{t}|_{t\in W_{1}})\neq\text{Law}(Y_{t}|_{t\in W_{2}})$
Our test statistic for this hypothesis is the Euclidean distance between the normalised signatures of the paths corresponding to the two windows:
$D=\|S(X_{W_{1}})-S(X_{W_{2}})\|_{2}$
where $S(X_{W_{i}})$ denotes the truncated signature of the normalised path in window $W_{i}$. We chose the Euclidean distance for its computational efficiency and empirical effectiveness, though other metrics such as the signature kernel distance could be explored in future work. Under the null hypothesis, we expect this distance to be small, reflecting the random variation of a single, consistent process. Under the alternative hypothesis, we expect the distance to be large, reflecting a genuine change in the underlying dynamics of the path. A sufficiently large value of this test statistic leads us to reject the null hypothesis and declare the presence of a change point at the boundary between $W_{1}$ and $W_{2}$.
4 Methodology: Signature-Based Change Point Detection
Our proposed methodology for detecting creative fatigue is based on a systematic algorithmic process designed to identify statistically significant changes in the geometric properties of a performance time series. This section details the algorithm, the selection of its key parameters, and its computational complexity.
4.1 Algorithm and Pseudocode
The core of our method is a four-step process that uses a sliding window to compare the geometric shape of a time series against itself.
First, we employ a sliding window approach. The algorithm moves a pair of adjacent, non-overlapping windows, $W_{1}$ and $W_{2}$, of a fixed size $w_{d}$ across the entire length of the time series. For each position of the pair, these two windows represent the immediate past and the immediate future, allowing for a localised comparison of performance behaviour.
Second, for each window, we compute its corresponding path signature. The time-series data within the window is first normalised to ensure that the signature calculation is not sensitive to the absolute scale of the time or metric axes. Time is scaled to the interval $[0,1]$, and the performance metric is min-max scaled, also to $[0,1]$. The normalised two-dimensional path is then passed to the library, which computes the path signature up to a pre-specified truncation depth $d$. This operation yields two signature vectors, $S(W_{1})$ and $S(W_{2})$, which serve as high-dimensional feature representations of the performance trajectory within each window.
Third, we quantify the difference in performance behaviour by calculating the Euclidean distance, $D=\|S(W_{1})-S(W_{2})\|$, between the two signature vectors. This distance provides a single, powerful measure of the dissimilarity in the shape of the performance paths in the two adjacent windows. A large distance implies a significant change in the underlying dynamics of the time series. To distinguish meaningful changes from random noise, we establish a statistical threshold for this distance. In our retrospective evaluation we compute the mean $\mu_{D}$ and standard deviation $\sigma_{D}$ of all the calculated distances across the time series for a given creative. A change point is then flagged at the start of window $W_{2}$ if its corresponding distance $D$ exceeds a sensitivity threshold defined as $\mu_{D}+k\sigma_{D}$, where $k$ is a user-defined parameter that controls the sensitivity of the detector. For online deployment, the same thresholding rule can be implemented using an initial calibration period or a rolling robust estimate of location and scale computed strictly from past distances.
Fourth, after all change points have been identified, the original time series is partitioned into a set of contiguous segments. For each segment, we perform a final trend classification. We fit a simple linear regression model to the data points within the segment and classify the trend as ’improving’, ’declining’, or ’stable’ based on the statistical significance and sign of the regression slope. This final step provides an interpretable label for the performance behaviour within each period of stability, making the results directly actionable for a marketing analyst. The complete process is summarised in Algorithm 1.
Algorithm 1 Signature-Based Change Point Detection
1:Input: Time series $Y=\{y_{1},...,y_{T}\}$, window size $w_{d}$, signature depth $d$, threshold multiplier $k$.
2:Output: List of change point dates CPCP, List of segment trends $T_{s}$.
3:Initialise $Distances\leftarrow[]$.
4:for $i\leftarrow 1$ to $T-2w_{d}+1$ do
5:  $W_{1}\leftarrow Y[i:i+w_{d}]$
6:  $W_{2}\leftarrow Y[i+w_{d}:i+2w_{d}]$
7:  $W_{1}^{\text{norm}}\leftarrow\text{Normalise}(W_{1})$ $\triangleright$ Scale time to [0,1], min-max scale metric to [0,1]
8:  $W_{2}^{\text{norm}}\leftarrow\text{Normalise}(W_{2})$
9:  $S_{1}\leftarrow\text{ComputeSignature}(W_{1}^{\text{norm}},d)$
10:  $S_{2}\leftarrow\text{ComputeSignature}(W_{2}^{\text{norm}},d)$
11:  $D_{i}\leftarrow\text{EuclideanDistance}(S_{1},S_{2})$
12:  Append $(Y[i+w_{d}].\text{date},D_{i})$ to $Distances$.
13:end for
[…]
20:for each SegmentSegment in SegmentsSegments do
21:  $Trend\leftarrow\text{ClassifyTrendViaRegression}(Segment,\alpha=0.05)$ $\triangleright$ $\alpha$ is significance level
22:  Append TrendTrend to $T_{s}$.
23:end for
24:return $CP,T_{s}$
4.2 Parameter Selection and Tuning
The performance of the signature-based change point detection algorithm is influenced by two key parameters: the window size $w_{d}$ and the sensitivity threshold multiplier $k$. The signature depth $d$ is a third parameter, but its impact is generally less critical provided it is large enough (e.g., $d\geq 3$) to capture non-trivial geometric information.
The window size $w_{d}$ determines the granularity of the analysis. A smaller window size will be more sensitive to short-term fluctuations and can detect changes more quickly, but it may also be more susceptible to noise and false positives. A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events. The optimal choice of $w_{d}$ depends on the nature of the time series being analysed. For daily data from typical digital advertising campaigns, we have found that a window size between 7 and 14 days provides a good balance between sensitivity and stability.
The sensitivity threshold $k$ controls the trade-off between false positives and false negatives. It determines how large the distance between two consecutive signatures must be to be considered a statistically significant change. A smaller value of $k$ will make the detector more sensitive, leading to more change points being identified, but potentially including more false positives. A larger value of $k$ will make the detector more conservative, reducing the number of false positives but increasing the risk of missing true changes. A value of $k$ between 1.5 and 2.5 has been found to work well in practice, providing a reasonable balance for typical performance marketing data. The optimal choice can be fine-tuned using a validation set if historical data with known change points is available.
4.3 Computational Complexity
The computational complexity of the algorithm is dominated by the sliding window loop. Let $T$ be the length of the time series and $w_{d}$ be the window size. The loop runs approximately $T-2w_{d}$ times. Inside the loop, the most expensive operation is the signature computation. For fixed path dimension $m=2$ and small truncation depth $d\in\{3,4\}$, signature computation is effectively linear in the number of time points per window. Formally, the number of truncated signature terms scales as $\sum_{k=0}^{d}m^{k}$, so computational cost grows quickly with $d$ but remains practical for small $d$. Since $w_{d}$ and $d$ are small constants in practice, the algorithm’s complexity scales linearly with the length of the time series, $\mathcal{O}(T)$. This linear scaling makes the methodology highly efficient and suitable for application to large-scale datasets containing many thousands of time series.
5 Quantifying Wastage: The Financial Impact of Fatigue
A key advantage of our methodology is its ability to translate a statistical observation into a tangible financial metric. While identifying a change point is analytically valuable, its true business utility is realised when we can quantify the cost of inaction. To this end, we introduce a framework for calculating ’wastage’, defined as the opportunity cost incurred by continuing to invest in a creative asset after its performance has begun to degrade.
The calculation first requires the establishment of a fair performance benchmark. This benchmark represents the level of performance the creative was capable of achieving during its peak. We establish this by analysing the segments identified by the change point detection algorithm. The benchmark CTR, denoted $CTR_{bench}$, is calculated as the average CTR over the most effective period of stable or improving performance. This ensures the creative is judged against its own demonstrated potential rather than an arbitrary external standard.
With the benchmark established, we can calculate the opportunity cost for any subsequent period where performance falls short. For each day $t$ after the benchmark period, we calculate the number of ’lost clicks’, which represents the additional clicks that would have been generated had the creative maintained its benchmark performance. To reduce sensitivity to day-to-day sampling noise, we compute lost clicks using the same noise-robust CTR estimate used in the onset definition below. This is given by the formula
$\text{Lost Clicks}_{t}=\max(0,CTR_{bench}-\widehat{p}_{t})\times\text{Impressions}_{t}$
where $\widehat{p}_{t}$ is the Beta–Binomial shrinkage estimate of CTR and $\text{Impressions}_{t}$ is the number of impressions on day $t$.
To translate these lost clicks into a monetary value, we introduce a cost benchmark, $CPC_{bench}$, which is typically the average cost-per-click during the same peak performance period. The daily wastage is then
$\text{Wastage}_{t}=\text{Lost Clicks}_{t}\times CPC_{bench}$
The total wastage for a creative is the sum of this daily wastage over the entire period of underperformance. This framework provides a clear and interpretable financial metric that directly links the statistical detection of fatigue to a business-relevant outcome. It allows marketing managers to not only identify underperforming assets but also to prioritise them based on the financial impact of their decline, leading to more efficient and data-driven budget allocation.
6 Experimental Validation
6.1 Synthetic Data Framework
To address the confidentiality constraints inherent in digital advertising data while enabling methodological exposition and evaluation, we use a synthetic panel dataset designed to mimic common characteristics of performance marketing time series: (i) high-frequency daily measurement, (ii) substantial day-to-day volatility in CTR, and (iii) heterogeneous trajectories across creatives.
6.1.1 Dataset structure and provenance
The synthetic dataset is included with the repository (data/synthetic/synthetic_data.csv). It contains 3,077 observations from 2025–01–01 to 2025–06–30 across 17 creatives spanning multiple campaigns. Each record contains impressions and clicks (allowing CTR to be treated as a noisy binomial rate), alongside auxiliary spend fields. Most creatives have near-daily coverage over the study window, but a subset are intentionally irregularly observed to reflect real-world missingness (e.g., pacing, pauses, or delivery gaps). The CSV also includes a pre-computed CTR column; throughout this paper we define $CTR_{t}:=C_{t}/I_{t}$ and use clicks and impressions for all modelling and evaluation.
6.1.2 Alignment with empirical magnitudes
The panel is calibrated to resemble plausible ranges reported in industry and marketing analytics contexts: impression volumes are typically in the thousands to tens of thousands per day, and CTR values concentrate in the low single-percent range (approximately 0.5%–3.5%), consistent with broad digital advertising benchmarks (IAB, 2023). The dataset is not intended to reproduce any specific platform or sector; rather, it provides a transparent setting for comparing detectors under a clearly specified operational ground truth.
Importantly, the dataset does not include injected changepoints or labelled fatigue onsets. We therefore define fatigue onset operationally in Section 6.4 and evaluate methods against that rule-based definition.
6.2 Comparative Method Analysis
We benchmark our signature-based approach against established methods from both marketing practice and statistical literature. Table 6.2 provides a comprehensive comparison:
[hb]
Comprehensive Comparison of Creative Fatigue Detection Methods Method Complexity Training Delay Robustness Interpretability Scalability Path Signature $O(n)$ No Tunable High Medium Excellent CUSUM $O(n)$ No 2-5 days Low High Excellent MA Crossover $O(n)$ No 3-7 days Very Low High Excellent ARIMA $O(n^{2})$ Yes 5-10 days Medium Low Poor Rolling Regression $O(n\,w_{reg})$ No 2-4 days Low High Good PELT1 $O(n)$ No 1-3 days Medium Medium Good Bayesian CPD2 $O(n^{2})$ Yes 0-2 days High Low Poor
  • 1(Killick et al., 2012)
    2(Adams and MacKay, 2007)
The signature-based approach offers several key advantages. Unlike machine learning methods, it requires no training and operates immediately on new creatives. Its responsiveness is tunable through the window size and threshold parameter, enabling a practical trade-off between early warning and alert burden. Its linear complexity enables scaling to portfolios containing thousands of creatives. Furthermore, the geometric robustness of the approach allows it to capture complex patterns beyond simple mean or variance shifts, making it particularly suited to the dynamic nature of advertising performance data.
6.3 Experimental Design
Our experimental validation follows a rigorous protocol:
  1. 1.
    Dataset: Use the synthetic panel dataset described in Section 6.1.
  2. 2.
    Operational Ground Truth: Define fatigue onset $\tau$ per creative via the noise-robust, benchmark-relative rule in Section 6.4.
  3. 3.
    Method Application: Apply the signature-based detector and two baseline heuristics (moving-average crossover; rolling regression) to each creative.
  4. 4.
[…]
  5. 5.
    Sensitivity Analysis: Vary detector parameters to characterise the trade-off between lead time and alert burden.
For the baseline methods, we use $CTR_{t}:=C_{t}/I_{t}$ and define windows over observed days (i.e., the last $k$ available observations for a creative) to accommodate missingness. Moving-average crossover uses trailing simple moving averages $MA_{7}(t)$ and $MA_{21}(t)$ computed over the last 7 and 21 observed days, respectively. An alert triggers on observed day $t$ if $MA_{7}(t)<MA_{21}(t)$ and this inequality holds for three consecutive observed days. Rolling regression fits an OLS model of $CTR$ on time over the last $w_{reg}=21$ observed days and triggers when the estimated slope is negative and statistically significant at $\alpha=0.05$ for three consecutive observed days.
The experimental framework ensures fair comparison by: - Using identical data for all methods - Using consistent pre-processing and smoothing assumptions in the operational ground truth - Defining a common alert date $\hat{\tau}$ for each method as specified in Section 7 - Reporting sensitivity analysis over detector parameters to characterise the early-warning versus alert-burden trade-off
6.4 Operational Definition of Fatigue Onset (Ground Truth)
Our empirical evaluation requires a reproducible definition of “fatigue onset” that is robust to the substantial day-to-day volatility inherent in digital advertising metrics. Because the dataset does not contain injected changepoints or labelled onset dates, we adopt an operational ground truth based on a sustained, economically meaningful deterioration in click-through rate (CTR) relative to a recent performance baseline.
Noise-robust CTR estimation.
Let $I_{t}$ and $C_{t}$ denote daily impressions and clicks for a given creative, and define the observed CTR as $CTR_{t}=C_{t}/I_{t}$. In finite samples, $CTR_{t}$ exhibits non-trivial measurement noise (particularly when $I_{t}$ fluctuates). To stabilise the series prior to defining fatigue onset, we use a simple Beta–Binomial shrinkage estimator:
$\widehat{p}_{t}=\frac{C_{t}+\alpha}{I_{t}+\alpha+\beta},$
where $(\alpha,\beta)$ represent a weakly informative prior over the underlying click probability. We parameterise the prior by a notional prior sample size $m_{0}$ and a global mean CTR $\bar{p}$, setting $\alpha=m_{0}\bar{p}$ and $\beta=m_{0}(1-\bar{p})$. In our experiments, we use $m_{0}=1000$ impressions and estimate $\bar{p}$ as the pooled mean CTR across all creatives. We then define the smoothed CTR series as a trailing moving average of length $w_{s}$:
$\widetilde{CTR}_{t}=\frac{1}{w_{s}}\sum_{j=0}^{w_{s}-1}\widehat{p}_{t-j},$
with $w_{s}=7$ to mitigate weekly noise while preserving responsiveness.
Recent-best baseline and onset criterion.
Fatigue is defined relative to a “recent best” baseline. For each day $t$, define
$Base_{t}=\max_{u\in[t-B,\;t-w_{s}]}\widetilde{CTR}_{u},$
where $B$ is the baseline lookback horizon (we use $B=21$ days). The relative degradation on day $t$ is then
$Drop_{t}=1-\frac{\widetilde{CTR}_{t}}{Base_{t}}.$
We declare that a creative has entered fatigue when the degradation exceeds a material threshold $\delta$ for a sustained period of $p$ consecutive days. Formally, the fatigue onset date $\tau$ is the earliest $t$ such that
$Drop_{t^{\prime}}\geq\delta\quad\text{for all}\quad t^{\prime}\in\{t,t+1,\dots,t+p-1\}.$
In the main specification, we set $\delta=0.15$ (a 15% sustained decline relative to recent-best performance) and $p=5$ days. Creatives that never satisfy this criterion are treated as having no fatigue onset within the observed horizon.
All quantities are evaluated only when the required history is available. For creatives with missing days, lookback windows are defined over the last $k$ observed days rather than calendar days. This onset definition is retrospective because it requires observing $p$ consecutive days; in an online setting, the earliest actionable alert corresponding to this definition would occur at day t+p1t+p-1.
Robustness checks.
To assess sensitivity to the operational definition, we report robustness results over $\delta\in\{0.10,0.15,0.20\}$ and $p\in\{3,5,7\}$, holding the smoothing parameters fixed. This ensures that comparative conclusions are not artefacts of a single threshold choice.
6.5 Case Study Visualisations
To demonstrate the practical application and interpretability of our methodology, we apply it to five synthetic datasets designed to mimic common real-world scenarios of creative performance decay. The resulting visualisations, shown in Figures 1 through 5, illustrate how the signature-based method identifies change points and classifies performance trends across a variety of fatigue patterns. Each plot displays the daily CTR, the detected change points as vertical dashed lines, and the classified trend for each segment.
Figure 1: Classic Wear-Out: A period of stable performance followed by a long, gradual decline.
Figure 2: Sharp Drop: A sudden, significant drop in performance after a stable period.
Figure 3: Fatigue and Recovery: A period of decline followed by a partial performance recovery.
Figure 4: Volatile Decline: A general downward trend characterised by high day-to-day volatility.
Figure 5: Multi-Stage Decline: A stepped decline with multiple distinct change points and periods of stability.
Figure 6: Non-Continuous Data (Sharp Drop): Demonstrates the method’s robustness to significant gaps in the data.
Figure 7: Non-Continuous Data (Gradual Decline): The signature method correctly identifies the change point despite irregular sampling.
These visualisations demonstrate the algorithm’s ability to adapt to different fatigue patterns. For a marketing analyst, these plots provide a clear and immediate narrative of a creative’s performance lifecycle, highlighting not just *that* performance has changed, but *when* and in what manner. Figures 6 and 7 specifically illustrate that signature-based methods can be applied to irregularly sampled trajectories. Window-based baselines can also be applied under missingness, but require an explicit choice of whether windows are defined over calendar days or observed days.
7 Results
We evaluate detection performance against the operational fatigue onset definition in Section 6.4. The dataset contains 17 creatives over 2025–01–01 to 2025–06–30, of which 6 meet the operational fatigue onset criterion under the main specification ($\delta=0.15$, $p=5$, $w_{s}=7$, $B=21$, $m_{0}=1000$). For each method, we define an alert date $\hat{\tau}$ as the first alert date for that creative. For the signature-based approach, $\hat{\tau}$ is the start date of the first segment classified as “declining” (with a fallback to the first detected change point). For baselines, $\hat{\tau}$ is the first day on which their heuristic trigger persists for three consecutive observed days.
[…]
Table 1: Lead time and alert burden under an operational ground truth definition
Method $N_{fat}$ $N_{hit}$ $N_{miss}$ $N_{burden}$ Median lead
Path Signature ($w_{d}=14$, $k=2.5$, depth=4) 6 6 0 11 28
MA Crossover (7/21, persist=3) 6 6 0 11 83
Rolling reg. (win=21, persist=3) 6 2 4 0 -3
7.1 Interpretation: early warning versus alert burden
Table 1 highlights a practical trade-off. Both the signature-based detector and the moving-average crossover provide early warning for all fatigue creatives under the operational definition, but they also generate alerts for all non-fatigue creatives in this dataset. By contrast, the rolling regression baseline is conservative (low burden) but misses most fatigue onsets and is, on average, late when it does trigger. Large positive lead times should be interpreted cautiously: extremely early alerts can reflect a mismatch between a heuristic trigger and the operational onset definition, rather than “better” detection.
7.2 Sensitivity analysis for the signature-based detector
We next vary the signature detector’s window size $w_{d}$ and threshold multiplier $k$ to characterise the lead-time versus alert-burden trade-off. Table 2 reports the number of fatigue creatives alerted (out of 6), burden (out of 11), and the median lead time over fatigue creatives.
Table 2: Sensitivity of signature-based detection to $w_{d}$ and $k$ (depth=4)
Window $w_{d}$ Threshold $k$ Alerts on fatigue Burden / Median lead (days)
7 1.5 6/6 11 / 66.5
7 2.0 6/6 11 / 48.5
7 3.0 2/6 7 / 35.0
14 1.5 6/6 11 / 92.0
14 2.5 6/6 11 / 28.0
14 3.0 4/6 8 / -5.5
7.2.1 Financial Impact Quantification
The true value of earlier detection becomes apparent when translated to operational units. The following calculation is a stylised illustration (not an empirical estimate from our synthetic panel). Given daily impressions $I$, benchmark cost-per-click $CPC_{bench}$, and an absolute CTR shortfall of $\Delta CTR$ relative to a benchmark period, the implied opportunity cost per day is
$\text{Opportunity Cost per day}=I\,\Delta CTR\,CPC_{bench}.$
This mapping is provided only to translate lead-time into operational units; it is not intended as a validated estimate of platform spending dynamics.
8 Practical Implementation and Managerial Implications
Beyond the theoretical and experimental validation, the practical utility of our methodology depends on its ease of implementation and the clarity of its implications for marketing decision-making.
8.1 Implementation Considerations
The signature-based detection algorithm is designed to be integrated into existing marketing analytics workflows. In terms of computational requirements, the algorithm exhibits linear complexity with respect to the length of the time series, as discussed in Section 4. This makes it highly scalable and efficient. A typical analysis for a single creative with a year’s worth of daily data can be completed in a fraction of a second on standard hardware.
The methodology can be deployed in either a batch or a real-time setting. In a batch processing mode, the analysis can be run periodically (e.g., daily or weekly) across an entire portfolio of active creatives, generating a summary report of assets that may be exhibiting fatigue. In a more advanced, real-time implementation, the algorithm could be integrated with a streaming data pipeline, allowing for the continuous monitoring of creative performance and the generation of automated alerts when a potential change point is detected. This would enable the fastest possible response to creative fatigue.
8.2 Managerial Implications
The primary implication of this work for marketing managers is the ability to move from a reactive, heuristic-based approach to creative management to a proactive, data-driven one. By providing a reliable and early warning system for creative fatigue, our methodology enables several key managerial actions.
First, it allows for more efficient budget allocation. Instead of continuing to spend on a creative that is no longer effective, managers can reallocate that budget to higher-performing assets or to the development of new creative. The wastage calculation provides a direct financial justification for these decisions.
Second, it provides a quantitative basis for the creative refresh cycle. The output of the analysis can be used to prioritise which creatives need to be replaced or updated most urgently, based on the magnitude of their performance decline and the associated financial wastage.
Third, it creates a more rigorous feedback loop for the creative development process itself. By analysing which types of creatives fatigue more quickly, marketing teams can gain insights into what resonates most with their audience over the long term, informing the development of more durable and effective advertising campaigns in the future.
9 Ethical Considerations and Reproducibility
9.1 Ethical Considerations
The methodology presented in this paper operates on aggregated, anonymised performance data. It does not require access to any personally identifiable information (PII) or individual user-level data, and as such, it aligns with modern data privacy principles such as those outlined in GDPR and other regulations. The analysis focuses on the performance of the creative asset itself, not on the behaviour of individual users. However, any application of performance measurement could indirectly lead to algorithmic bias. For example, if the system consistently flags creatives targeted at a specific demographic as fatiguing more quickly, it could lead to a reduction in advertising shown to that group. It is therefore important that practitioners use this tool as one input into a broader decision-making process, and remain mindful of the potential for such second-order effects.
10 Conclusion
In this paper, we have introduced a novel methodology for identifying and quantifying advertising creative fatigue by applying the mathematical framework of path signatures. Our findings from a synthetic panel evaluation indicate that this signature-based approach can provide early warnings under an explicit operational onset definition, while highlighting a practical trade-off between early warning and alert burden. By treating the performance time series as a geometric path and using its signature as a rich feature descriptor, our method can detect changes in the underlying dynamics of creative performance that simpler models may miss. The subsequent quantification of this fatigue in terms of opportunity cost provides a bridge from statistical outputs to operational prioritisation.
The practical implications of this methodology for marketing practitioners are significant. The automated and scalable nature of the algorithm allows for the continuous monitoring of a large portfolio of creative assets, something that is infeasible to do manually. This enables marketing and media teams to move from a reactive to a proactive stance on creative management. Rather than waiting for performance to decline substantially, they can be alerted to the onset of fatigue in near real-time. The wastage metric provides a clear, data-driven rationale for when to retire an underperforming creative and reallocate its budget, thereby improving the overall efficiency and return on investment of advertising campaigns.
Appendix A A Deeper Dive into Rough Path Theory
Rough path theory provides a mathematical framework for describing the interactions between complex, oscillating systems. At its core is the concept of the path signature, which serves as a faithful, non-parametric summary of a path. For a continuous path $X_{t}$ in $\mathbb{R}^{m}$, its signature $S(X)$ is an infinite sequence of tensors, living in the tensor algebra space over $\mathbb{R}^{m}$.
The first level of the signature, $S^{1}(X)$, is a vector representing the total displacement of the path.
$S^{1}(X)_{i}=\int_{0}^{1}dX_{t}^{i}$
The second level, $S^{2}(X)$, is a matrix whose entries are iterated integrals.
$S^{2}(X)_{ij}=\int_{0}^{1}\int_{0}^{t}dX_{s}^{i}dX_{t}^{j}$
These second-order terms capture the area swept out by the path in the $ij$-plane. Higher-order terms capture more complex geometric properties.
A key result from rough path theory, Chen’s theorem, states that the signature uniquely determines the path up to tree-like deformations. This means that if two paths have the same signature, they are essentially equivalent from a geometric perspective. This uniqueness property is what makes the signature such a powerful feature set for describing path-like data. In practice, we use a truncated signature, which provides a finite-dimensional approximation that is sufficient for most machine learning applications. Our library provides an efficient and numerically stable implementation for computing these truncated signatures.
Appendix B Algorithm Parameters
The performance of the signature-based change point detection algorithm is influenced by two key parameters: the window size $w$ and the sensitivity threshold $k$.
The window size $w$ determines the granularity of the analysis. A smaller window size will be more sensitive to short-term fluctuations and can detect changes more quickly, but it may also be more susceptible to noise and false positives. A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events. The optimal choice of $w$ depends on the nature of the time series being analysed. For daily data from typical digital advertising campaigns, we have found that a window size between 7 and 14 days provides a good balance between sensitivity and stability.
The sensitivity threshold $k$ controls the trade-off between false positives and false negatives. It determines how large the distance between two consecutive signatures must be to be considered a statistically significant change. A smaller value of $k$ will make the detector more sensitive, leading to more change points being identified, but potentially including more false positives. A larger value of $k$ will make the detector more conservative, reducing the number of false positives but increasing the risk of missing true changes. A value of $k$ between 1.5 and 2.5 has been found to work well in practice, providing a reasonable balance for typical performance marketing data. The optimal choice can be fine-tuned using a validation set if historical data with known change points is available.
Appendix C Mathematical Foundations and Theoretical Analysis
This section provides informal theoretical support for why signature distances are sensitive to changes in the underlying trajectory. For readability, we present proof sketches rather than fully detailed proofs. We assume paths are continuous and of bounded variation (or, more generally, finite $p$-variation) so that truncated signatures are well-defined. To avoid notational ambiguity, we use $m$ to denote the path dimension and $d$ to denote the signature truncation depth.
C.1 Signature distance sensitivity
Theorem 1 (Signature sensitivity to path changes (proof sketch)).
Let $X,Y:[0,T]\to\mathbb{R}^{2}$ be two continuous paths representing performance trajectories, with signatures $S(X)$ and $S(Y)$ truncated at depth $d$. If there exists a time interval $t_{1},t_{2}]\subset[0,T]$ where the paths exhibit different geometric properties (trend, volatility, or curvature), then:
$\|S(X)_{[t_{1},t_{2}]}-S(Y)_{[t_{1},t_{2}]}\|\geq C\cdot\sup_{t\in[t_{1},t_{2}]}\|X(t)-Y(t)\|,$
where $C>0$ depends on regularity assumptions and the truncation depth $d$.
Proof.
The signature captures iterated integrals of the path increments. For the first level:
$S^{1}(X)_{[t_{1},t_{2}]}-S^{1}(Y)_{[t_{1},t_{2}]}=\int_{t_{1}}^{t_{2}}(dX_{t}-dY_{t})=X(t_{2})-X(t_{1})-(Y(t_{2})-Y(t_{1}))$
This directly captures the difference in total displacement. For the second level, which captures signed area:
$S^{2}(X)_{[t_{1},t_{2}]}^{ij}-S^{2}(Y)_{[t_{1},t_{2}]}^{ij}=\int_{t_{1}}^{t_{2}}\int_{t_{1}}^{t}(dX_{s}^{i}dX_{t}^{j}-dY_{s}^{i}dY_{t}^{j})$
This term is sensitive to differences in the path’s oscillatory behaviour and volatility. By the Chen-Strichartz formula, the signature norm satisfies:
$\|S(X)-S(Y)\|\geq\|S^{1}(X)-S^{1}(Y)\|\geq\|X(T)-X(0)-(Y(T)-Y(0))\|$
For paths with different geometric properties (e.g., one declining, one stable), the higher-order signature terms amplify these differences, providing the stated lower bound. A rigorous statement requires explicit norms and assumptions (e.g., bounded $p$-variation) and follows from continuity estimates for truncated signatures. ∎
C.2 Theoretical Analysis of Detection Power
Theorem 2 (Detection power (heuristic bound)).
Consider a time series $\{Y_{t}\}$ with a change point at $\tau$. Let the process have different distributions before and after $\tau$:
$Y_{t}\sim\begin{cases}\mathcal{F}_{1}&\text{if }t<\tau\\ \mathcal{F}_{2}&\text{if }t\geq\tau\end{cases}$
For window size $w$ and threshold $\theta=\mu_{D}+k\sigma_{D}$, the detection power (probability of correctly identifying the change) is:
$\text{Power}=P(D>\theta|H_{1})\geq 1-\exp\left(-\frac{w\cdot\text{KL}(\mathcal{F}_{1}||\mathcal{F}_{2})}{2\sigma^{2}}\right)$
where $\text{KL}(\cdot||\cdot)$ is the Kullback-Leibler divergence and $\sigma^{2}$ is the variance of the signature distance under the null hypothesis.
Proof.
Under $H_{1}$ (change present), windows $W_{1}$ and $W_{2}$ straddling the change point contain samples from different distributions. The signature distance $D$ between these windows can be bounded using the data processing inequality:
$D^{2}=\|S(X_{W_{1}})-S(X_{W_{2}})\|^{2}\geq w\cdot d_{\text{TV}}(\mathcal{F}_{1},\mathcal{F}_{2})^{2}$
where $d_{\text{TV}}$ is the total variation distance.
By Pinsker’s inequality:
$d_{\text{TV}}(\mathcal{F}_{1},\mathcal{F}_{2})\geq\sqrt{\frac{1}{2}\text{KL}(\mathcal{F}_{1}||\mathcal{F}_{2})}$
Under $H_{0}$, the signature distances concentrate around their mean; a complete derivation depends on dependence assumptions in the time series and on the distribution of the signature estimator. Using concentration-style reasoning:
$P(D>\theta|H_{1})\geq 1-\exp\left(-\frac{(\mathbb{E}[D|H_{1}]-\theta)^{2}}{2\sigma^{2}}\right)$
Substituting lower bounds on $\mathbb{E}[D|H_{1}]$ yields the stated qualitative dependence. The intent is to highlight that detection improves with larger window sizes and with greater separation between pre- and post-change regimes. ∎
C.3 Convergence Properties as Window Size Increases
Theorem 3 (Asymptotic consistency (informal)).
Let $w$ be the window size and $n$ be the total number of observations. As $w\to\infty$ with $w=o(n)$, the signature-based change point estimator $\hat{\tau}_{w}$ converges to the true change point $\tau$ in probability:
$P(|\hat{\tau}_{w}-\tau|>\epsilon)\to 0\quad\text{as }w\to\infty$
for any $\epsilon>0$.
Proof.
Define the localised signature distance process:
$D_{t}(w)=\|S(X_{[t-w,t]})-S(X_{[t,t+w]})\|$
Under regularity conditions (e.g., bounded $p$-variation and appropriate mixing), this process satisfies:
  1. 1.
    For $t$ far from $\tau$: $\mathbb{E}[D_{t}(w)]=O(w^{-1/2})$ (by CLT for signatures)
  2. 2.
    For $|t-\tau|<w$: $\mathbb{E}[D_{t}(w)]\geq c\cdot\min(|t-\tau|/w,1)$ for some $c>0$
The change point estimator is:
$\hat{\tau}_{w}=\arg\max_{w<t<n-w}D_{t}(w)$
By the continuous mapping theorem and the fact that $D_{t}(w)$ has a unique maximum near $\tau$ (with high probability), we have:
$\hat{\tau}_{w}\to\tau\quad\text{in probability}$
The rate of convergence depends on additional assumptions and is stated here for intuition only. Larger window sizes provide more stable estimation at the cost of reduced temporal resolution. ∎
C.4 Computational Complexity Analysis
Lemma 1 (Signature computation complexity (order of growth)).
For a path of length $w$ in $\mathbb{R}^{m}$ and truncation depth $d$, the number of truncated signature terms grows as $\sum_{k=1}^{d}m^{k}$. A naive computation scales approximately as:
$\mathcal{O}(w\cdot m^{d}\cdot d).$
In practice, log-signature representations reduce constant factors substantially; for fixed $m=2$ and small $d$ (as used in our experiments), the overall detection pipeline is effectively linear in the number of time points.
Proof.
The dominant cost is evaluating iterated integrals (or equivalent algebraic updates) over $w$ increments. Since the number of terms grows exponentially in $d$, practical implementations fix $d$ (e.g., $d\in\{3,4\}$) and treat it as a small constant. ∎
C.5 Robustness to Noise and Outliers
Theorem 4 (Noise robustness (informal)).
Let $X(t)$ be the true performance path and $\tilde{X}(t)=X(t)+\epsilon(t)$ be the observed path with noise $\epsilon(t)$ satisfying $\mathbb{E}[\epsilon(t)]=0$ and $\text{Var}(\epsilon(t))=\sigma^{2}_{\epsilon}$. Then the expected signature distance satisfies:
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq C\cdot w\cdot\sigma^{2}_{\epsilon}\cdot\left(1+O(w^{-1/2})\right),$
where $C$ depends on the truncation depth $d$ and regularity of $X$.
Proof.
The signature is a continuous functional of the path. By the rough path stability theorem, small perturbations in the path lead to controlled changes in the signature. For additive noise:
$S(\tilde{X})-S(X)=S(X+\epsilon)-S(X)=\sum_{k=1}^{d}\int\cdots\int d\epsilon_{t_{1}}\otimes\cdots\otimes dX_{t_{k}}+O(\|\epsilon\|^{2})$
Taking expectations and using the independence of noise:
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]=\sum_{k=1}^{d}\mathbb{E}\left[\left\|\int\cdots\int d\epsilon_{t_{1}}\otimes dX_{t_{2}}\otimes\cdots\right\|^{2}\right]$
By Itô isometry and the bounded variation of $X$:
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq\sigma^{2}_{\epsilon}\cdot w\cdot\sum_{k=1}^{d}\frac{\text{Var}(X)^{k-1}}{k!}$
The sum converges to a finite constant under bounded-variation assumptions, giving the stated qualitative dependence. This shows that the signature-based method has linear sensitivity to noise variance, with the normalisation step providing additional robustness. ∎